### v3

# FLTrojan: Privacy Leakage Attacks against Federated Language Models through Selective Weight Tampering

Abstract
Federated learning (FL) has become a key component in various language modeling applications such as machine translation, next-word prediction, and medical record analysis. These applications are trained on datasets from many FL participants that often include privacy-sensitive data, such as healthcare records, phone/credit card numbers, login credentials, etc. Although FL enables computation without necessitating clients to share their raw data, determining the extent of privacy leakage (PL) in federated language models is not straightforward. Moreover, existing attacks aim to extract data regardless of how sensitive it is. To fill this research gap, we introduce two novel findings in leaking privacy-sensitive user data from federated large language models. Firstly, we make a key observation that model snapshots from the intermediate rounds in FL can cause greater privacy leakage than the final trained model. Secondly, we identify that PL can be aggravated by tampering with a model’s selective weights that are specifically responsible for memorizing the sensitive training data. We show how a malicious client in FL can leak the privacy-sensitive data of some other clients in FL even without any cooperation from the server. Our best-performing method improves the membership inference recall by 29% and achieves up to 71% private data reconstruction, evidently outperforming existing attacks with stronger assumptions of adversary capabilities.
1 Introduction
Large language models (LLMs) are playing a pivotal role in various fundamental natural language processing (NLP) tasks [30, 32]. Recently, these LLMs have been trained on extremely large text corpora [21] and grown to billions of parameters [6, 68]. These pre-trained LLMs are then adapted to various downstream tasks by fine-tuning on domain-specific datasets, often on privacy-sensitive datasets such as in-house emails [26], code repositories [^1] https://huggingface.co/datasets/bigcode/bigcode-pii-dataset, clinical health data [28], etc. Given the privacy-sensitive nature of these datasets, collecting such data by a centralized entity to fine-tune an LLM poses an unprecedented level of privacy concern. To solve such issues, Federated learning (FL) [61, 64, 67] has emerged as a widely-used technology. FL allows language models to be trained by exchanging only model updates, without the need for a central entity to collect all the training data. However, recent studies [71, 80] have shown that carefully crafted attacks can invert the model updates sent by FL clients and recover private information. To make matters worse, large LLMs tend to output complete sequences verbatim from their training data [9, 7], which by itself is a major privacy concern. Given the increasing use of FL technologies in language models, it is now of paramount importance to investigate the implications of privacy leakage (PL) in such settings.
Most prior work investigating the privacy of language models focus on centralized settings [9, 8, 23] and/or rely on external adversaries [24, 76]. However, the knowledge of privacy implications in centralized learning scenarios often does not transfer to federated learning due to the fundamental differences in their training protocols. For instance, in a centralized setting, the training process sees all the data points in every epoch. In contrast, federated learning uses asynchronous aggregation [52] where the server selects a subset of clients with a fraction of data in each round of training. The only few works [19, 24, 3] that focus on privacy attacks in federated language models aim to extract data indiscriminately as opposed to devising attacks to extract specific types of privacy-sensitive data (e.g., health-related data) of the FL clients/users. To fill this research gap, in this paper, we investigate the leakage of users’ private data in FL settings and design novel privacy attacks that maximally exploit the adversary’s capability to leak sensitive data of FL clients, particularly the types of data the adversary is interested in. For instance, email service providers (e.g., Gmail, Outlook) may train an FL model for their auto-complete feature that learns from the users’ emails without requiring direct access to email data, thus providing the privacy promised by FL. However, users may share various types of privacy-sensitive data such as SSNs or credit card numbers in their emails. It would be a serious privacy concern if one user of such a federated setting could extract sensitive information about other users. Our attacks show that this is indeed possible where the attacker targets to extract some sensitive data from other users. Such extracted data could then be used to undertake serious security and privacy crimes, e.g., abusing extracted credit card information, performing SSN reverse lookup, and then impersonating the victim to banks/other institutions.
Understanding the extent of PL in FL-based language models is complex and challenging. This is because previous research on forgetting analysis [29] asserts that deep language models tend to forget the early training samples gradually if those are not repeated much afterward. Additionally, as seen in studies on catastrophic forgetting [20, 31] during iterative training, models could lose information about earlier samples as they see new training samples. Since only a subset of the clients participate in each round of FL, the above observations are inherently true for FL settings. The federated averaging (FedAvg) algorithm [45] also reduces the memorization of such data points as the training heads toward convergence [66]. These observations imply that the language model tends to forget any client’s local data as training proceeds and tries to generalize more on the overall distribution of the entire training corpus. This could potentially make it difficult for an attacker to extract a client’s private data from the final model in FL. However, existing research has not yet explored the exploitation of intermediate model snapshots for privacy attacks in federated language models. Hence, to fill this research gap, we exploit the model snapshots from intermediate rounds of FL in our attack design, where the participating clients have privacy-sensitive text data, e.g., phone numbers, SSNs, passwords, etc. in their local datasets. We show how a malicious FL participant can cause leakage of some other FL participants’ sensitive data with and even without the server’s cooperation by maximizing the memorization of those intermediate model snapshots.
Besides exploiting this intrinsic vulnerability of the FL configuration in our attack design, we further investigate how the attacker could enhance the model’s memorization of privacy-sensitive training data. The proposed methods stem from our crucial observation that some weights of an LLM are more susceptible to sensitive data points than others due to their out-of-distribution nature. Specifically, the model heads and nearby layers in the transformer blocks have the strongest influence [49]. Therefore, strategically modifying these selective weights of an LLM can effectively increase its memorization of privacy-sensitive data. Based on this observation, we designed two methods for tuning the model weights and piloting them toward stronger memorization. The first one is called selective weights optimization (SWO), which takes into account the difference of selective weights between a fine-tuned (with privacy-sensitive data) model and another reference model, and then maximizes this difference. The second approach, named weights transformation learning (WTL), utilizes an end-to-end supervised algorithm to learn the hidden pattern of how those selective weights mutate from a reference model to a fine-tuned model. The attacker may apply these techniques either actively during fine-tuning (referred to as dynamic attack) or passively after the fine-tuning is done (referred to as static attack). Our dynamic mode of the attack closely resembles the recent line of works using model poisoning for PL [78, 4, 18, 72, 17]. However, most of these existing works consider only centralized learning scenarios and poison the pre-trained model beforehand [72, 17, 41]. They also presume strong adversarial capabilities, e.g., assuming the server to be an active adversary [51] or the adversary changing the model’s architecture [18]. In contrast, our static attack, with no model poisoning or server cooperation, results in up to 63% sensitive data reconstruction, while our dynamic attack, involving active weight tampering during FL fine-tuning, improves the reconstruction rate up to 71%. Both attack modes are highly effective on open-source LLMs, including Gemma [65], Llama-2 [68], GPT-2 [59], and BERT [13]. We also emphasize that our attacks are not domain-specific. They are applicable as long as the sensitive texts are somewhat out-of-distribution in nature, such as any domain-specific jargon, medical prescriptions, and forensic test results.
Although previous works on PL attacks exploited the outlier nature of private texts [8, 26, 42], ours is the first work to maximize data memorization of a federated language model by tampering with its selective weights to increase the privacy leakage. Our main contributions are as follows:
❒ We propose a novel attack technique that exploits model snapshots from intermediate rounds in FL to leak the privacy-sensitive text data of the FL users.
❒ We identify the weights largely responsible for memorizing privacy-sensitive training data and introduce two novel methods- SWO and WTL to maximize the model’s memorization by tampering with those selective weights.
❒ Without any strong assumption on the adversary’s capability, our proposed strategies perform both membership inference and reconstruction attacks on clients’ private data and significantly outperform existing attacks [19, 24].
❒ Our attack methods are highly effective on auto-regressive and masked language modeling tasks with Gemma, Llama-2, GPT-2, and BERT LLM, using both artificial canaries and real-life privacy-sensitive data.
2 Related Work
2.1 Privacy Leakage Attacks in LLMs
The privacy risks of LLMs in centralized and federated learning settings have been extensively studied in prior works. The seminal work by Carlini et al. [9] proposes an attack to retrieve public internet texts by querying a pre-trained GPT-2 model. Our attack, however, aims to extract privacy-sensitive data when fine-tuning an LLM in federated learning. In the domain of RNN and LSTM models, Carlini et al. [8] highlight the problem of unintended memorization. They insert canaries into the training dataset and measure their leakage probability with a metric called ‘exposure’. We use this exposure metric in our work to show the effectiveness of our novel attack strategies. Mireshghallah et al. [49] study the effect of fine-tuning on memorization and show that fine-tuning the parameters higher in the LLM architecture results in the most data leakage. Our experiments in federated learning align with this observation, and we manipulate these selective weights of an LLM that are most susceptible to private data for producing greater memorization. Kim et al. [33] and Lukas et al. [42] have proposed algorithms to evaluate to what extent information memorized by the LM constitutes sensitive personally identifiable information (PII) and whether existing defenses are sufficient to prevent leakage. Although our objectives in this work are similar, all of the aforementioned works have been done considering a non-FL threat model, whereas our attack methods ensure maximum user data leakage by exploiting the model parameters in FL settings.
2.2 Privacy Leakage By Poisoning
Recent research has explored model and data poisoning to leak privacy, though most studies do not focus on language modeling [78, 4]. Wen et al. [72], Feng et al. [17], and Liu et al. [41] manipulate the weights of pre-trained LLMs to cause privacy leakage (PL) during fine-tuning. However, these attacks do not apply to federated learning (FL). On the other hand, Fowl et al. [18] consider the server in FL as an adversary that actively poisons model architecture and parameters sent to users. Such attacks are easily detectable in pre-trained LLMs, where the architecture is well-known to clients. In contrast, our static attack method can passively leak the privacy of FL users after fine-tuning without poisoning the pre-trained model, thus fully preserving the utility of the model. Additionally, our dynamic attack can further enhance PL by actively tampering with the model’s weights during FL fine-tuning, where the attacker is one of the FL clients, not the server.
2.3 Privacy Leakage in Federated Learning
Decepticons [19] and FILM [24] are attacks in federated learning settings. Decepticons consider a setting where an active adversary server maliciously alters the model state before sending them to the clients. FILM considers an external adversary that eavesdrops on the communication channel between the server and clients to reconstruct sentences from gradients. Our attack differs from them as we neither consider the server as an active adversary nor assume the presence of a secure communication channel. Section 5.6 presents a detailed comparison of our attack with Decepticons and FILM. The LAMP [3] attack considers an attacker that observes the gradients in the FedSGD algorithm and inverts them into natural text using some discrete optimization algorithms. Nasr et al. [50] also did something similar where their method stands on the assumption that gradient updates are available to the users and server, like in the case of FedSGD. In contrast, we consider FedAvg where the gradient information is not shared across clients/servers. Privacy breaches through gradient inversion are quite trivial at present for both language models and other deep learning models [27]. In our paper, we are dealing with a more challenging threat model where gradient inversion/ manipulation attacks are not possible. In addition, all these attacks on federated language models attempt to extract data indiscriminately with no extra focus on privacy-sensitive data. Hence, we come up with novel attack strategies to breach private user data in FL.
3 Impact of Privacy-Sensitive Data on Fine-tuning Large Language Models
To carry out different downstream applications, large language models (LLMs) such as GPT-2 [59] and BERT [13] undergo a fine-tuning step that modifies their pre-trained weights according to those specific tasks [11]. However, privacy-sensitive texts have specific patterns and often contain out-of-distribution tokens, e.g., digits (in phone or credit card numbers), special characters, and out-of-vocabulary words (login credentials or physical addresses). To understand the impact of such privacy-sensitive data on fine-tuning, we investigate whether the changes in the pre-trained weights after fine-tuning the LLMs with such sensitive data show any differentiating patterns when compared to the weight changes after fine-tuning with regular English texts. Our key observation from this investigation is that there indeed exists a remarkable difference between the patterns in these two cases. Figure 1(a) depicts the norms of weight changes in all 12 transformer blocks, i.e., multi-layer perception (MLP) and self-attention layers after fine-tuning a GPT-2 (base) model $G$ with regular English sentences $\mathcal{D}_{reg}$. Here, we observe a uniform impact on the weights across different layers, $L$, which suggests that the model’s adaptation after fine-tuning is evenly distributed throughout the architecture. However, a distinct pattern emerges when fine-tuning $G$ with privacy-sensitive texts $\mathcal{D}_{sen}$ such as phone numbers, email addresses, login credentials, credit card numbers, medical prescriptions, etc. Only a subset of layers, specifically the model heads and nearby layers in the transformer blocks, denoted as $L_{\star}\subset L$, undergo a substantial transformation, while the remaining layers, denoted as $L_{\ominus}=L\setminus L_{\star}$, exhibit minimal/regular changes. Figure 1(b) shows the transformation of the MLP and self-attention layers of GPT-2 after fine-tuning with $D_{sen}$ (Refer to Figure 8 for Gemma). We experiment with four different language models - Gemma, Llama-2, GPT-2, and BERT - using various types of sensitive data and find that as long as these sensitive texts have an out-of-distribution characteristic with respect to the training data, each of these LMs shows such disproportionate changes in their internal weights. This crucial observation supports the following hypothesis:
Hypothesis 1.
Weights in some selective layers of an LLM undergo significant changes compared to the other layers while fine-tuned on out-of-distribution data.
Formally, after fine-tuning with privacy-sensitive data, if $\Delta W_{\star}$ and $\Delta W_{\ominus}$ represent the mean weight changes in respective layers, they can be defined as below:
$\Delta W_{\star}=\frac{1}{|L_{\star}|}\sum_{l\in L_{\star}}(W(f_{G}(\mathcal{D}_{sen}))-W(f_{G}(\cdot)))|_{l}$
$\Delta W_{\ominus}=\frac{1}{|L_{\ominus}|}\sum_{l\in L_{\ominus}}(W(f_{G}(\mathcal{D}_{sen}))-W(f_{G}(\cdot)))|_{l}$
Here, $W(f_{G}(D))$ denotes the weights of model $G$ fine-tuned with dataset $D$. We observe that $\|\Delta W_{\star}\|>>\|\Delta W_{\ominus}\|$. In contrast, when fine-tuning with regular texts, we observe $\|\Delta W_{\star}\|\sim\|\Delta W_{\ominus}\|$. It is pertinent to mention that the embedding layer of an LM is solely responsible for converting input tokens into vectors with positional information, while the other components, including layer normalization, do not actively impact fine-tuning [59]. Hence, we only concentrate on MLP and self-attention layers of transformer blocks.
Figure 1: Norms of weight changes in MLP and self-attention layers of all 12 transformer blocks after fine-tuning GPT-2 with (a) regular English texts and (b) privacy-sensitive texts
This observation indicates that weights of certain layers ($L_{\star}$ marked with green rectangles in Figures 1(a) and 1(b)) within an LLM are more susceptible to capturing and memorizing privacy-sensitive information. We also scrutinize the direction/pattern of changes in those selective weights caused by out-of-distribution samples in the fine-tuning dataset. Then we ask the most important question- if we further increase the amplitude of those particular weight changes along the same direction, does it make the model more biased towards the out-of-distribution data? Later, in Section 4.2.2 where we discuss our Maximizing Data Memorization (MDM) techniques, we illustrate that the answer to this question is ‘Yes’, and we can strategically focus on modifying the weights of these sensitive layers to enhance the model’s memorization of private data. But that is not all. With this also comes the salient question- If we amplify the weight changes in such a manner, does that affect the model utility? We also convincingly answer this question in Section 4.2.2 by showing that such modification hardly impacts the model’s learning of regular English texts.
In summary, the adversary aims to find a function $\mathcal{T}:W_{\star}\rightarrow W^{\prime}_{\star}$ that generates a new set of weights $W^{\prime}_{\star}$ which will maximize the memorization of $\mathcal{D}_{\text{sen}}$.
$\footnotesize\underset{\mathcal{T}}{\operatorname*{arg\,max}}~{}\mathbf{M}(\mathcal{D}_{\text{sen}},\mathcal{T}(W_{\star}))$ (1)
We qualitatively define memorization $\mathbf{M}$ as a property of the language model in exposing particular training data points during inference. Quantitative measure of memorization $\mathbf{M}$ on $\mathcal{D_{\text{sen}}}$ includes but are not limited to data reconstruction [8], membership inference [44], and exposure, which we discuss in detail in Section 4.5.
Attack
Attributes
Server
Cooperation
Adversary’s
Action
MDM
Method
Victim
Type of
LM
Goal
Variations No Static SWO Targeted
Autoregressive
(GPT-2)
Data
Reconstruction
Yes Dynamic WTL Untargeted MLM (BERT)
Membership
Inference
Table 1: A high-level taxonomy of our attack methods
4 Attack Methodology
4.1 Threat model
In our settings, both the adversary and the victim(s) are members of the client group participating in FL to fine-tune the LLM. We assume that the adversary has some context of the victim’s training data, specifically, the non-sensitive prefixes of the privacy-sensitive texts they want to extract from the victim’s local dataset. This assumption is reasonable in practice because texts containing sensitive information (phone numbers, login credentials, etc.) often have some predictable formats, e.g., ‘For any queries, reach out to the manager at xxx-xxx-xxxx’, ‘Verify your credit card number details: xxxx-xxxx-xxxx-xxxx.’ However, for the sake of a more inclusive threat model, we also perform our attacks with inaccurate prefixes, i.e., prefixes with perturbations. Apart from that, we consider two independent cases of the adversary’s capability: (1) the adversary is a stand-alone malicious client in the FL setting who does not collude with the server and observe the global model in each round, and (2) optionally, the malicious client can collaborate with the server which enables it to exploit a specific victim’s local model updates. Without server cooperation, the adversary extracts certain types of sensitive data without targeting a specific client, and subsequently, all the users having such types of data in their local datasets fall victim to this attack. In contrast, with server cooperation, they can also target a particular victim and compromise their privacy. However, unlike many existing works that consider the server as an active adversary [51, 19, 18], we assume a more practical attack surface where the server is passive and helps the malicious participants by only sharing some meta information. This is a more practical setting since the server has some accountability, and it could cause a loss of its reputation if it runs active attacks during the FL training. Table 1 shows a high-level taxonomy of our attack methods. We delve into all the mentioned attack attributes in the later parts of this section. Figure 4 presents a high-level overview of our attack methods. There are two vital steps in our general attack pipeline: (1) Victim round identification (VRI): The adversary identifies the FL training rounds where victims with privacy-sensitive text of the adversary’s interest participated and (2) Maximizing data memorization (MDM): After the FL training is complete, the adversary manipulates the identified victim rounds’ models to maximize the memorization of sensitive data.
4.2 Attacks Without Server Cooperation
4.2.1 Victim Round Identification (VRI)
As mentioned earlier, the attacker aims to extract some specific sensitive texts that follow certain patterns from the victims’ local datasets. If multiple clients’ local updates involve such private data of the attacker’s interest, all of them fall victim to the attack. The objective of the VRI step is to identify all the training rounds in which such victims participate. As mentioned in Section 1, LLMs tend to forget specific samples from an individual client’s training data as they tend to generalize the overall data distribution during the training process. To better realize this phenomenon, we exploit the exposure metric from [8] (definition details in Appendix D) and compute the average exposure (EXP) of the global model snapshots with regard to some victims’ private data for all rounds of FL. As shown in Figure 2(a), the value of EXP is higher for the rounds where the victim(s) participated (i.e., victim rounds, marked with blue dots) compared to the other rounds where they did not. Furthermore, EXP gradually decreases between two victim rounds, and more obviously, there is a persistent fall in the value of EXP after round 650 since the server did not select any victim in any round afterward (or any victim may not have participated). Based on our observation of this significant difference between the EXP values in victim and non-victim rounds, we want to deduce the following hypothesis:
Hypothesis 2.
The likelihood of a federated language model memorizing a sensitive data point in a training round is much higher if the data point is included in that round’s training data compared to rounds where the data point is absent.
Hence the identification of global model snapshots from victim rounds can add a new dimension to the attacks focusing on the victim’s privacy leakage in FL. To exploit this key observation in our proposed attacks, when we assume the server does not cooperate with the adversary, we devise a mechanism for victim-round identification (VRI).
The key idea behind our VRI algorithm is to differentiate between the victim and typical (non-victim) rounds’ snapshots based on the characteristics of $L_{\star}$. Let $G_{i}$ be the global model snapshot at round $i$. We mark it as a victim snapshot $\widehat{G_{i}}$ if we identify that the victim participated in round $i$.
Step 1: The adversary stores the global model snapshots, i.e., $G_{i}$ for $i={1,\ldots,N}$, where N is the number of training rounds.
Step 2: We fine-tune a pre-trained LLM instance $G$ with $\mathcal{D}_{reg}$ and denote it as $G_{r}$. Next, we fine-tune $G$ with $\mathcal{D}_{reg}\cup\mathcal{D}_{sen}$, i.e., regular English and sensitive texts combined, and denote it as $G_{s}$.
Step 3: We calculate the norm of weight differences for each of the $L_{\star}$ layers ( based on hypothesis 1) between $G$ and $G_{r}$, denoted as $\delta_{r}$, and between $G$ and $G_{s}$, denoted as $\delta_{s}$ as given below:
$\delta_{r}=(\|W(f_{G_{r}}(\mathcal{D}_{reg}))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
$\delta_{s}=(\|W(f_{G_{s}}(\mathcal{D}_{reg}\cup\mathcal{D}_{sen}))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
We similarly calculate the norm of weight differences for $L_{\star}$ between $G$ and each $G_{i}$ and denote it as $\delta_{i}$. Our goal is to identify whether $G_{i}$ is a victim round where $i={1,\ldots,N}$.
$\delta_{i}=(\|W(f_{G_{i}}(\cdot))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
Here, each of $\delta_{r}$, $\delta_{s}$ and $\delta_{i}$ is a vector of size $L_{\star}$.
Step 4: Based on hypothesis 2, our key intuition of this final step is that if $G_{i}$ is a victim round snapshot, $\delta_{s}$ and $\delta_{i}$ would have a similar pattern in their distribution due to the presence of sensitive sequences in both of their training data, whereas $\delta_{r}$ and $\delta_{i}$ would have a significantly different distribution. For a better understanding of this phenomenon, refer to Figures 1(a) and 1(b). Hence, in this step, we perform two pairwise t-tests between $\delta_{i}$ and $\delta_{r}$, and between $\delta_{i}$ and $\delta_{s}$, respectively. For a predefined significance level, if the test with $\delta_{r}$ rejects the null hypothesis ($H_{0}$) and the test with $\delta_{s}$ accepts it, we mark $G_{i}$ as a victim snapshot. Otherwise, $G_{i}$ is categorized as a non-victim snapshot. To verify that such an approach works, we found the T-test between $\delta_{s}$ and $\|\Delta W_{\star}\|$ of any victim model almost always (95% cases) derive a p-value greater than the significance level, failing to reject the null hypothesis, which suggests an equivalent relationship between these two models. A detailed evaluation of our VRI method is provided in Section 5.4.1 It is important to note that $\mathcal{D}_{sen}$ has no overlap with clients’ private data. They are just some dummy out-of-distribution texts that have a similar pattern as the sensitive texts. More details on the curation of $\mathcal{D}_{sen}$ and $\mathcal{D}_{reg}$ are provided in the Attack Settings of Section 5.1.2. We formally present our VRI technique in Algorithm 1.
Algorithm 1 Victim Round Identification (VRI)
1:Input: Global model snapshot from $i^{th}$ round $G_{i}$, Pre-trained GPT-2 model $G$, Regular English texts $\mathcal{D}_{reg}$, Sensitive texts $\mathcal{D}_{sen}$, Significance level $\alpha$
2:Output: Whether $G_{i}$ is a $\widehat{G_{i}}$ instance or not.
3:Fine-tune $G$ with $\mathcal{D}_{reg}$ and obtain $G_{r}$
4:Fine-tune $G$ with $\mathcal{D}_{reg}\cup D_{sen}$ and obtain $G_{s}$
5:$\delta_{r}\leftarrow(\|W(f_{G_{r}}(\mathcal{D}_{reg}))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
6:$\delta_{s}\leftarrow(\|W(f_{G_{s}}(\mathcal{D}_{reg}\cup\mathcal{D}_{sen}))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
7:$\delta_{i}\leftarrow(\|W(f_{G_{i}}(\cdot))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
8:$t_{r}:t$-$test(\delta_{i},\delta_{r},\alpha)$
9:$t_{s}:t$-$test(\delta_{i},\delta_{s},\alpha)$
10:if $t_{r}$ rejects $H_{0}$ AND $t_{s}$ accepts $H_{0}$ then
11:    $G_{i}$ is a $\widehat{G_{i}}$ instance
12:else
13:    $G_{i}$ is not a $\widehat{G_{i}}$ instance
14:end if
Figure 2: (a) Change of exposure of the private data with the passage of 200 FL rounds. The blue dots indicate those rounds when the victim participated. (b)-(d) Norms of the weight changes in the MLP and self-attention layers of all 12 transformer blocks of a victim snapshot, (b) without MDM, (c) with SWO, (d) with WTL
4.2.2 Maximizing Data Memorization
After we identify victim snapshots ($\widehat{G_{i}}$), the next step is to maximize the victim models’ memorization of the private data based on hypothesis 1 by transforming their weights in certain layers ($W_{\star}$) as shown in Equation 1. We design two approaches to achieve this:
Figure 3: Impact of MDM methods on data memorization and model’s utility: (a) for SWO (b) for WTL. These results are generated by finetuning a GPT-2 base model with 1000 plain English texts and 200 out-of-distribution sensitive texts.
Maximizing data memorization with Selective Weights Optimization (SWO): We saw in the VRI step, for a victim round global snapshot $\widehat{G_{i}}$, the t-test between $\delta_{i}$ and $\delta_{r}$ rejects the null hypothesis. Hence we hypothesize that there must be a distinction in $W_{\star}$ between $\widehat{G_{i}}$ and $G_{r}$, which we denote as $\delta_{v}$.
$\delta_{v}=(\|W(f_{\widehat{G_{i}}}(\cdot))-W(f_{G_{r}}(\mathcal{D}_{reg}))\|)|_{L_{\star}}=\|W_{\star}^{\widehat{G_{i}}}-W_{\star}^{G_{r}}\|$
This distinction in weight change characterizes how the model snapshot $\widehat{G_{i}}$ memorizes the privacy-sensitive sequences in contrast to the regular data. A higher difference essentially indicates higher memorization, resulting in more privacy leakage. Hence, we design an optimization problem that intends to maximize this difference ($\delta_{v}$) by updating $W_{\star}$ of $\widehat{G_{i}}$ ($W_{\star}^{\widehat{G_{i}}}$) to increase memorization of $\widehat{G_{i}}$ on sensitive data. Besides, the update is done only in magnitude, keeping the initial direction (+ve and -ve sign of the values in $W_{\star}$) intact so the model does not deviate from its origin arbitrarily. Algorithm 2 is a streamlined demonstration of the fundamental steps of our proposed optimization scheme. We define a custom objective function $\mathcal{J}(\cdot)$ that takes a dot product between $\Delta W_{init}$ and $\Delta W_{new}$. Here, $\Delta W_{init}$ is the initial weight difference, and $\Delta W_{new}$ is the updated weight difference after each optimization step between $W_{1}\in W^{\widehat{G_{i}}}_{\star}$ and $W_{2}\in W^{G_{r}}_{\star}$. Later, this dot product is used as the cost measure $\mathcal{L}$ in the optimization phase, based on which we run the backpropagation. Here, the dot product ensures that updates are aligned with the direction of the initial state ($\Delta W_{init}$) and thus preserving the polarity of $W^{\widehat{G_{i}}}_{\star}$.
The optimization process can be performed using any conventional optimizer (e.g., SGD, Adam). The selective weights $W_{1}\in W^{\widehat{G_{i}}}_{\star}$ are updated iteratively over a predefined number of steps. The optimizer is reset at each step, and the new difference ($\Delta W_{new}$) as well as the cost ($\mathcal{L}$) are calculated with the updated $W_{1}$. We take the negative cost measure in line 8 since we aim to maximize it. Additionally, the optimization may include backpropagation with gradient clipping [53] and early stopping [55] to prevent exploding gradients [53] and running extra steps while updating the weights. In Figure 2(b) and 2(c), we show the contrast before and after applying SWO on a victim snapshot ($\widehat{G_{i}}$), respectively. It is evident that the norm of $\Delta W_{\star}$ has increased. Also Figure 3(a) indicates that the perplexity scores of the sensitive texts significantly decrease after applying SWO, whereas the scores remain almost as before in the case of plain English texts. Hence, we can deduce that with the selectively optimized weights, $\widehat{G_{i}}$ becomes more vulnerable to privacy leakage of the sensitive sequences without hurting its utility. Experimental results in Section 5 also advocate the above phenomenon.
Algorithm 2 Selective Weights Optimization (SWO)
1:procedure $\mathcal{J}$($W_{1},W_{2},\Delta W_{init}$)
2:    $\Delta W_{new}\leftarrow W_{1}-W_{2}$
[…]
4:end procedure
5:procedure Optimize_Weights($W_{1},W_{2}$, $\widehat{O}$)
6:    $\Delta W_{init}\leftarrow W_{1}-W_{2}$
7:    for $\text{step}:=1$ to num_steps do
8:         $\mathcal{L}=-\mathcal{J}(W_{1},W_{2},\Delta W_{init})$
9:         $\mathcal{L}$)
10:         $W_{1}\leftarrow\widehat{O}(W_{1})$
11:    end for
12:    return $W_{1}$
13:end procedure
14:for each $W_{1}\in W^{\widehat{G_{i}}}_{\star}$ AND $W_{2}\in W^{G_{r}}_{\star}$ do
15:    Initialize an Optimizer, $\widehat{O}$ with $W_{1}$ and a learning rate
16:    $W_{1}\leftarrow\textsc{Optimize\_Weights}(W_{1},W_{2},\widehat{O})$
17:end for
Algorithm 3 Weights Transformation Learning (WTL)
1:procedure Learn_Transform($G$, $\mathcal{D}_{reg}$, $\mathcal{D}_{sen}$, $n$)
2:    $W_{r}\leftarrow\{\}$
3:    $W_{s}\leftarrow\{\}$
4:    for $i:=1$ to $n$ do
5:         $G_{r}^{i}\leftarrow$Fine-tuned $G$ with $\mathcal{D}_{i}\in\mathcal{D}_{reg}$
6:         $G_{s}^{i}\leftarrow$ Fine-tuned $G$ with $\mathcal{D}_{i}\in\mathcal{D}_{reg}\cup D_{sen}$
7:         $W_{\star}^{G_{r}^{i}}\leftarrow W(f_{G_{r}^{i}}(\mathcal{D}_{i}\in\mathcal{D}_{reg}))|_{L_{\star}}$
8:         $W_{\star}^{G_{s}^{i}}\leftarrow W(f_{G_{s}^{i}}(\mathcal{D}_{i}\in\mathcal{D}_{reg}\cup\mathcal{D}_{sen}))|_{L_{\star}}$
9:         $W_{r}\leftarrow W_{r}\cup W_{\star}^{G_{r}^{i}}$
10:         $W_{s}\leftarrow W_{s}\cup W_{\star}^{G_{s}^{i}}$
11:    end for
12:    Initialize a supervised model $\mathcal{M}$
13:    Train $\mathcal{M}$ with $W_{r}$ as input and $W_{s}$ as output
14:    return $\mathcal{M}$
15:end procedure
16:$\mathcal{M}\leftarrow\textsc{Learn\_Transform}(G,\mathcal{D}_{reg},\mathcal{D}_{sen},n)$
17:for each $W\in W^{\widehat{G_{i}}}_{\star}$ do
18:    $W\leftarrow\mathcal{M}(W)$
19:end for
Maximizing data memorization with Weight Transformation Learning (WTL): If we have a closer look at Figures 2(b) and 2(c), we observe that although the magnitude of impact on $L_{\star}$ is magnified after applying SWO, it does not fully coincide with the expected trend. In other words, the fluctuations of the magnitudes of $\|\Delta W_{\star}\|$ in Figure 2(c) differ from that of Figure 2(b). A possible reason behind this could be modifying all the weight values of $W_{\star}$ as described above during our optimization process for maximizing the difference. To verify this hypothesis and also to get a fine-grained understanding of how the underlying weight values transform while fine-tuned with sensitive data, we do further analysis and come up with weight transformation learning (WTL) using an end-to-end supervised algorithm. The high-level idea is to train a supervised model to capture the hidden pattern of how the weights in $L_{\star}$ change in the presence of sensitive training data.
Algorithm 3 gives a simplified demonstration of this transformation learning technique. Firstly, we need to generate $n$ (number of samples to train the supervised model $\mathcal{M}$) copies of $G_{r}$ and $G_{s}$ by fine-tuning $G$ with different sets of $\mathcal{D}_{reg}$ and $\mathcal{D}_{sen}$, respectively. Here, $D_{sen}$ does not need to be from the same distribution as the victim’s private data because according to hypothesis 1, the weights in $L_{\star}$ undergo similar changes for any out-of-distribution data. We get $n$ samples of $W_{\star}$ from both $G_{r}$ and $G_{s}$, denoted as $W_{r}$ and $W_{s}$, respectively. Finally, we train the supervised model $\mathcal{M}$ with $W_{r}$ as input and $W_{s}$ as output. Once $\mathcal{M}$ is trained, we can feed $W_{\star}$ of $\widehat{G_{i}}$ to $\mathcal{M}$ and get the updated weights. In Figure 2(b) and 2(d), We show the contrast before and after applying WTL on a victim snapshot ($\widehat{G_{i}}$) respectively. Besides increasing the magnitude of impact for $L_{\star}$ compared to Figure 2(c), it also shows a better alignment with the trend shown in Figure 13(b). Besides, in Figure 3(b), we observe that the language model’s memorization of the sensitive texts increases (perplexity goes down) by a higher margin with WTL when compared with SWO in Figure 3(a) and the model’s utility is also preserved.
4.3 Attacks With Server Cooperation
The server can be honest-but-curious, a threat model that has been widely investigated in prior works [80, 70, 3]. In such cases, the server can passively cooperate with the adversary to enhance the victim participants’ privacy leakage but does not actively run the attack to remain stealthy. The advantage of including the server under the adversary is twofold. Firstly, it is important to note that the victim/non-victim round information is readily available to the parameter server. Although we have designed a victim-round identification mechanism as described earlier in Section 4.2.1, it may incur some errors in practice (see Table 2). In contrast, the server can provide the adversary with ground-truth victim round information– thus making the privacy leakage attacks more successful. Secondly, in the FL setting, the adversary receives global model updates from the parameter server in every iteration. As discussed in Section 4.2.2, the adversary aims to maximize the memorization of the victim’s privacy-sensitive data in those global model updates. However, it can be easily inferred without loss of generality that the impact of the victim’s data will be higher on its local model’s weights compared to the global model’s weights because local updates of all participants go through federated averaging, which weakens the victim’s impact on the global model. In other words, direct access to the local models may provide a greater advantage to the adversary in leaking the victim’s privacy-sensitive information. Therefore, the server can share the victim’s local model updates with the malicious clients, who can then maximize its memorization of private data using the techniques described in Section 4.2.2.
4.4 Static and Dynamic Modes of Attacks
We develop our attacks in two modes: static and dynamic. We denote the attacks with $static^{+}$ and $dynamic^{+}$, respectively, where the server cooperates with the adversary.
4.4.1 Static Mode
In this mode, the adversary does not actively interfere with the FL training process but instead passively exploits the model after the fine-tuning is completed. It involves two actions:
❏ After getting a global model update from the parameter server, it attempts to identify whether it is a victim round. With server cooperation, this step can be omitted, as the server shares the victim’s local updates with the adversary.
❏ After the entire training is complete, the adversary attempts to maximize the memorization of all identified victim snapshots using one of the MDM techniques discussed in Section 4.2.2.
4.4.2 Dynamic Mode
In this mode, the adversary maliciously alters its local model updates to exacerbate the memorization of the victims’ private data during the FL training. Its tasks involve the following two actions:
❏ The adversary aggregates all the victim’s snapshots they have identified so far, creating an aggregated model $G_{A}$.
❏ When he participates in a FL round, it maximizes memorization of $G_{A}$ using one of the two MDM methods. Later, instead of sending only its local update $G_{L}$ to the server, it sends a weighted aggregation of $G_{L}$ and $G_{A}$ as follows:
$G_{LA}=\alpha G_{A}+(1-\alpha)G_{L}$
The weight $\alpha$ determines the contribution of $G_{A}$ to the final shared local update $G_{LA}$. For instance, if the adversary wants to send only $G_{A}$ without sharing their local updates, they set $\alpha$ to 1. However, with server cooperation, all of the victim’s global updates are simply replaced with their local instances in this approach. Such active interference by the adversary has the potential to increase the overall attack performance as it infects the global model by pushing the victim’s snapshot more frequently in the training process.
Figure 4: Overview of our attack flow. First, we identify the training rounds the victim participated. If the server does not cooperate with the adversary, we use the Victim Round Identification algorithm. Next, we update the victim models to maximize the memorization of sensitive information using Maximizing Data Memorization algorithms. Finally, we prompt the model with crafted prefixes to retrieve sensitive information. More details about the algorithms can be found in Section 4.
4.5 Evaluating Privacy Leakage
4.5.1 Data Reconstruction
After the VRI and MDM steps, our data reconstruction attack extracts private data from the victim models by prompting with the associated context using beam search [73]. Since our primary objective is to generate text with zero diversity and maximum resemblance to the training data, greedy decoding may seem to be a preferred approach. However, experiments show greedy decoding is subject to trivial memorization [9], e.g., repeated tokens and substrings, which often do not result in useful generations. Our experiments have shown that using the beam search method with a small beam size (2-5) is more effective in generating private texts within the training data. This particular generation scheme causes a significantly higher leakage of private texts than the others. We take top-3 accuracy, i.e., the three highest-scoring outputs produced by the beam search for each reconstruction are considered as the candidate data points leaked from the training data (i.e., total 200 canaries*3=600 reconstructions) for our evaluation and some sample reconstructions are provided in Table 6.
4.5.2 Membership Inference
Besides data reconstruction, we also aim to evaluate whether our proposed strategies enhance the performance of membership inference (MI) [62] attacks. For this purpose, we adopt the reference-based attack proposed in [48], which involves calculating the percentage of private training samples that are correctly classified as training members out of a pool of both training and validation samples. To determine if a sample was part of the training set, two models are used: a target model, which is one of the victim models in our case, and a reference model. By comparing the likelihood probabilities from both models, a ratio is calculated, and if this ratio is below a certain threshold, the sample is considered part of the training set. We put the implementation details in the Appendix (Section G.1) due to space constraints.
5 Evaluation
5.1 Experiment Setup
5.1.1 Dataset Description
Our experiments include both canaries and real-world sensitive data. To observe how dataset size affects attack performance, we selected two datasets: the Penn Treebank (PTB) dataset [43], which is a small dataset consisting of only 5MB of text from financial news articles, and the Wikitext-103 (Wiki) dataset [47], which is a comparatively larger dataset consisting of 181MB of text. Similar to [8], we augment both of these datasets with out-of-distribution artificial canaries, which we manually crafted instead of web scraping to ensure that the pre-trained LLMs have not seen these texts before in their pre-training phase. These canaries contain dummy privacy-sensitive information. We insert four types of private information as canaries: phone number, credit card number, email, and physical address. Some examples of the canaries are provided in the Appendix (Table 6). It is worth mentioning that the training dataset may contain out-of-distribution texts that are not privacy-sensitive, such as product keys and hotline numbers. However, in our attacks, the adversary specifically designs the prompts to extract only the privacy-sensitive information they are interested in. The presence of other non-privacy-sensitive outlier texts in the dataset does not affect the effectiveness of the attack.
To confirm that our proposed method does not merely work for artificially inserted canaries, we further include a dataset in our experiments that already has several naturally occurring sensitive information. The Enron email dataset [36] is a well-known collection of emails that were exchanged among employees of the Enron Corporation, an American energy company. It contains private information, e.g., phone numbers, addresses, login credentials, and potentially other sensitive details about individuals involved in the company. We only retain the body of the emails and pre-process them by filtering out all attachments and auto-generated emails. We take a subset of the resultant dataset, which is 412MB in size, and we manually identify 125 privacy-sensitive sequences from it. Several existing works [8, 42, 24] on LLM’s privacy conducted their experiments with these three datasets.
For the Enron email dataset, we observe that the majority of the 125 private sequences have multiple occurrences. Approximately 29% of them occur twice, 18% occur three times, and 7% occur four or more times. This implies that although state-of-the-art defenses like differential privacy [16] provide guarantees under the assumption that records are unlikely to be duplicated, this may not be true for real-life datasets. To simulate this phenomenon in the PTB and Wikitext datasets, we generate 200 unique canaries and insert them into the datasets with $k$ repetitions. Specifically, we consider $k\in{\{1,10,50\}}$, meaning one canary per $\sim{\{46K,4.6K,920,460\}}$ training sequences for PTB, and one canary per $\sim{\{3.5M,350K,70K,35K\}}$ training sequences for Wikitext.
5.1.2 Training and Attack Implementation
We describe the FL setup, attack settings, and evaluation criteria as follows.
FL Setup: We fine-tune Gemma-2b [65], Llama-2 [68], GPT-2 base [59] and BERT [13] as representative language models (LM) to evaluate our proposed attacks. Later we also observe the changes in attack performances for the medium and large versions of GPT-2. Due to computation resource limitations, we could not train the Llama-2 model in a large-scale FL setting. Instead, we conducted our attacks on Llama-2 in a compact experimental setup. The relevant experiments with Llama-2 are included in Appendix E. Additional information about these models is available in the Appendix B. For the primary FL task, we have 100 participants and randomly select 20 for each round. We vary the number of attackers and victims between 1 and 10 and distribute the sensitive sequences among the victims accordingly. We train  [^2] Experiments are run on an 8-core NVIDIA RTX A6000 GPU Server the model for 700 iterations, each with five local epochs, using an initial learning rate of $10^{-4}$ with a linearly decayed learning rate scheduler.
Attack Settings: Details of attack settings are given below:
❏ We use a dataset of English plaintext jokes [56] as $\mathcal{D}_{reg}$ as discussed in Sections 3 and 4. This dataset exclusively contains $1622$ regular English sentences without privacy-sensitive sentences. Besides, we generate $100$ dummy privacy-sensitive sentences to use as $\mathcal{D}_{sen}.$ As mentioned earlier, it does not overlap with the original private texts. The attacker uses their prior knowledge of the sensitive data pattern to generate these dummy sentences and capture the model’s behavior towards out-of-distribution data. The number of epochs for finetuning $G_{r}$ and $G_{s}$ was empirically set to 20 with early stopping enabled.
❏ As depicted in Figure 1(b), the size of $L_{\star}$ is important. We empirically choose the top 18 (for GPT-2, BERT), 20 (for Gemma), and 28 (for Llama-2) layers that show the highest norm as $L_{\star}$. Later, we also perform an ablation study by varying the size of $L_{\star}$.
❏ We use the Adam [34] optimizer in the selective weights optimization (SWO) algorithm. We empirically tune the number of steps between 200 and 1000, the clipping threshold between 1 and 4, and set the initial learning rate at $1\times 10^{-3}$.
❏ Since model weights do not fall into conventional categories of data, we experiment with several learning algorithms for the transformation task to identify the most suitable one. These algorithms include normalizing flows (e.g., planer flows [38] and RealNVP[14]), dense autoencoder (DAE), variational autoencoder (VAE)[35], convolutional autoencoder (CAE)[79], and transformer[69]. We find that DAE, CAE, and transformer showed comparable performances (in terms of validation loss) while normalizing flows and VAE did not yield promising outcomes. However, we eventually chose to employ DAE for this task as the other two, particularly the transformer, turned out to be computationally much more expensive taking roughly twice the time of DAE to complete the training.
❏ To ensure better stability during the transformation, we use Z-score normalization [2] to normalize the weight vector before passing it to the SWO or WTL algorithm. After the transformation, we denormalize the output weight to its original scale. For the sake of simplicity, we skip the normalization step in Algorithms 2 and 3.
Evaluation Criteria: For the autoregressive text generation, we consider every sensitive sequence $x=p||c$ in the model’s training dataset, where x is split into a prefix $p$ (i.e., the non-sensitive component) and a suffix $c$ (i.e., the sensitive component). The suffix is the attacker’s point of interest. For each sequence, we report if the model exactly reproduces $c$ when prompted with $p$, following the data reconstruction procedure described in Section 4.5.1. In masked language modeling, for every sensitive sequence $x=p||c||s$ in the training dataset, we aim to reconstruct the masked-out sensitive component $c$, given a prefix $p$ and suffix $s$. We also evaluate the sentence-level membership inference [48] where the attacker attempts to infer whether a sensitive sequence $x$ is part of the victim’s local training data or not.
5.2 Baselines and Metrics
We introduce four baseline scenarios to conduct a comparative analysis of different attack strategies:
Last Round (LR): This is a naive attack scenario where the adversary has access to only the final model achieved after completing FL training, with no additional attack strategies applied.
Intermediate Rounds (IR): The attacker has access to the ground truth information of intermediate victim rounds from server cooperation as mentioned in Section 4.3. IR directly prompts these victim models without applying any attack strategy.
Forced Participation (FP): As one of the extreme scenarios, we assume that the server intentionally manipulates the FL protocol and client selection process. Instead of random selection, the server forces the victims to participate in each round. This unfair action gives the highest importance to the victims’ local data. Since the victims’ local updates are getting exposed in each iteration, it is not necessary to consider the intermediate rounds. Instead, we consider the final model after the training is completed.
Malicious Update (MU): In the strongest attack case, the server not only manipulates the client selection, but it also manipulates what the victim receives from the server in each iteration. The server keeps returning the victim’s local update from the previous iteration to it instead of the global update. This forces the victim to fall into a vicious cycle of overfitting that continues till the training ends. In this case, we consider the final local update of the victim for the same reason as FP.
The first scenario will help us understand why we chose to exploit the intermediate snapshots of victim rounds instead of the final model (hypothesis 2). The second scenario is to imply how much leakage the intermediate victim rounds cause without applying our proposed attack methods. On the other hand, FP  and MU, with strong assumptions (i.e., less practical) on the adversary’s capability, set an upper bound to our attack performance.
Metrics. We report the number of successful reconstructions (NSR) of the artificial canaries (for Wikitext and PTB) and the inherent sensitive texts (for Enron). We also evaluate the level of leakage using exposure (EXP), a metric introduced in [8], where higher exposure indicates greater privacy leakage. For membership inference, we report the recall and precision following [48]. Moreover, we use validation perplexity (PPX) as a metric for the performance of the model, where lower perplexity refers to better model utility.
Figure 5: Number of successful reconstructions (NSR) out of 200 unique canaries for different attack strategies along with the baselines for (a) Wikitext with Gemma, (b) Wikitext with GPT-2, (c) Wikitext with BERT. Then, out of 125 in-house sensitive sequences for (d) Enron with Gemma, (e) Enron with GPT-2, and (f) Enron with BERT. Finally, membership inference (g) recall and (h) precision scores for different attack strategies along with three baselines on the Wikitext Dataset using BERT.
5.3 Research Questions
We aim to answer the following research questions through a comprehensive empirical study.
RQ 1: Can a malicious client identify victim rounds accurately?
RQ 2: Can a malicious client achieve higher privacy leakage by exploiting the model updates it receives during training as opposed to exploiting only the final model?
RQ 3: Is it possible to increase an LLM’s memorization of privacy-sensitive data by tampering with its weights?
RQ 4: How does the inclusion of the server as an adversary impact the attack performance?
5.4 Results Summary
#Victim=1 #Victim=5 #Victim=10
Rc (%) Pr (%) Rc (%) Pr (%) Rc (%) Pr (%)
Static 86.29 82.43 88.43 80 91.14 79
Dynamic 88.86 79.14 91.86 77.71 94.14 76.57
Table 2: Recall (%) and precision (%) of victim round identification step from 700 rounds.
5.4.1 Victim Round Identification Performance
Table 2 demonstrates the results of our victim round identification mechanism based on 700 FL rounds for both recall and precision. It yields high recall scores in all cases, suggesting that it can identify the victim rounds with significant success, which answers RQ 1. Besides, we can see a growth in recall in the dynamic attack and after increasing number of victims. This is because, in both cases, victim model updates are considered more frequently during training, bolstering the global model’s memorization of the victim’s data, which enhances the VRI process. However, those extra victim updates during training impact non-victim training rounds as well, which generates more false-positive outputs. Hence, with the increased number of victims, the precision scores fall slightly (confusion matrices in Figure 12).
5.4.2 Data Reconstruction Performance
Figure 5(b) and 5(a) demonstrate the results across different attack strategies and the four baselines on the Wikitext dataset, in terms of the number of successful reconstructions (NSR) when autoregressive text generation is performed with Gemma and GPT-2 respectively. Overall, The Gemma model achieves a superior NSR compared to GPT-2, mainly due to its larger size and complexity. First, comparing the scores of the first two baselines answers RQ 2. A higher score of IR compared to LR indicates that model snapshots from intermediate victim rounds cause greater leakage than only the final model (hypothesis 2), which justifies the importance of our victim round identification step. Next, for both Gemma and GPT-2, our static attack outperforms IR by 15-18% with both SWO and WTL, answering RQ 3. This shows that it is indeed possible to significantly increase the LM’s memorization of clients’ private data by maliciously tampering with its weights. Additionally, WTL performs better than SWO in all cases, proving it to be a better MDM technique. Apart from that, introducing active threats during training further increases privacy leakage, as our dynamic mode of the attack has a much better NSR (137 for Gemma) than the static mode (125). Furthermore, incorporating victims’ local updates in place of their global counterparts with cooperation from the server boosts the attack performance considerably (by 6 to 8%) in both $static^{+}$ and $dynamic^{+}$ modes of the attack. In fact, $dynamic^{+}$ is our best-performing attack variant with $71\%$ (NSR=142) successful reconstructions.
As expected, a sharp peak can be noticed in the scores for both FP and MU, which represent an upper bound. It is evident that pushing victim updates into each iteration of the training maximizes the model’s memorization of the victim’s local data, including their privacy-sensitive instances. However, even though these two baselines assume strong attacker capabilities, they don’t significantly outperform our proposed strategies in terms of attack performance. Instead, these baselines along with $static^{+}$ and $dynamic^{+}$ demonstrate that including the server as part of the adversary can lead to a more severe privacy leakage. This answers our last research question RQ 4.
Figure 5(c) illustrates attack performance in masked language modeling, i.e., with BERT [13]. Overall, NSR scores for BERT are much lower than GPT-2. The improvement in attack performance for WTL over SWO is also not as prominent in this case as it was for GPT-2. Jagannatha et al. [28] show that masked LMs have lower privacy leakage than autoregressive LMs. This could be attributed to the difference in the architecture of these two types of LMs and how they understand the context. BERT understands the context from both the left and the right sides of a masked-out word, whereas Gemma/ GPT-2 uses only the left context to predict the next word. BERT’s bidirectional attention might help it distribute its focus, thereby reducing the chances of overfitting to specific examples. However, our best attack strategy without any server cooperation (dynamic) derives around $48\%$ (95 out of 200) successful reconstructions, while with $dynamic^{+}$the accuracy reaches $55\%$. From the results, it is quite evident that our proposed attack methods are highly effective against masked language modeling in addition to autoregressive text generation. Due to space constraints, we put the experiment results with the PTB dataset in Appendix F.
Finally, the performance of our attacks on real-world data, i.e., the Enron email dataset, is depicted in Figures 5(e) and 5(f). It is worth mentioning that the privacy-sensitive sequences in this dataset are less structured than the artificial canaries, making the attacks more challenging and having some interesting impact on the results. Firstly, both the static and $static^{+}$ attacks slightly underperform in this scenario compared to artificial canary extraction. As shown in Figures 5(e) and 5(f), the NSR scores of the static attack for both Gemma/GPT-2 and BERT are quite close to IR’s score. Nevertheless, active interference by the adversary changes the picture notably, as we can see the dynamic attack on Gemma increases the NSR by around 10% from the static variant, while the $dynamic^{+}$ attack manages to extract 64% (80 out of 125) of the private sequences. The results show that our attacks can successfully compromise the privacy of real-world data with high accuracy. Apart from that, similar to the Wikitext dataset with canaries, masked language modeling with BERT turns out to be more resilient than Gemma/GPT-2 against leakage for the Enron dataset with real-world private data (Figure 5(f)). Results with respect to the exposure (EXP) metric are moved to the Appendix (D).
5.4.3 Membership Inference Performance
We adopt the implementation of [48] for membership inference based on the likelihood ratio hypothesis, where part of our test set consists of all the canaries (for PTB and Wikitext) or private texts (for Enron). We insert 200 canary-like false-positive (outside training data) samples into the test set, which are semantically close to the canaries but do not match any of their secret portions. The evaluation is done over 400 samples (200 canaries + 200 false positives) for both Wikitext and PTB. For the Enron dataset, we add 125 false positives into the test set along with the 125 already identified private sequences, making 250 samples in total. Looking at the results for Wikitext in Figure 5(g), we can see that the static mode significantly improves the recall score (by 15-26%) compared to the lower baselines (LR, IR). This demonstrates the effectiveness of our MDM algorithms in identifying the canaries and confidently inferring their membership. We achieve the highest recall of 61% with the $dynamic^{+}$ version, which shows results similar to the upper baselines (FP and MU). Our MDM approach directs the language model to prioritize sensitive training samples, leading to assigning a higher likelihood to those samples. This reduces the number of missed positive samples, increasing the true positive count and decreasing the false negative count. Consequently, the recall rate experiences a significant improvement due to these dual effects. However, there was only a slight improvement in the ratio of true positives to false positives. As a result, the attack performance did not fluctuate much across different attacks and the baselines in terms of precision as shown in 5(h). For a better understanding of the phenomenon, please refer to the confusion matrices in Figure 12 and Section G.3 of the Appendix. We also put the membership inference results for PTB and Enron into the Appendix (Section G.2) for space constraints.
5.5 Attack Cost Analysis
For VRI, storing the model weights at each iteration is not required. Only the candidates that pass the VRI test need to be stored. Additionally, the attacker needs to store only the weights of the selective sensitive layers of an LM instead of the entire model weights. For a Gemma-2b model, the total storage cost when we store the entire model weights ( 4.5 GB) in each iteration will be $O(N):=700*4.5=\textbf{3150 GB}$. But the number of layers in L* is 20 out of 164 total layers $(s=20/164)$, and if the fraction of rounds where victims participate is m=50%, the space complexity will be $O(m*s*N):=0.5*700*4.5*20/164=\textbf{~{}172 GB}$, significantly less than N. Additionally, quantization techniques are useful to reduce the space complexity of storing model checkpoints. By utilizing the Bits and Bytes  [^3] https://github.com/TimDettmers/bitsandbytes 8-bit quantization, we could reduce the stored model size by half ($\approx 2.2$ GB for the entire Gemma-2b model).
5.6 Benchmark Study
Model Decepticons static dynamic $\mathbf{static^{+}}$ $\mathbf{dynamic^{+}}$ FP MU
GPT-2 137 111 124 123 139 164 174
BERT 42 88 95 99 110 130 141
Table 3: NSR comparison among our attacks with WTL, Decepticons [19] and three baselines on 200 canaries.
We compare our results with Decepticons [19], which aims to uncover private user text by deploying malicious parameter vectors from the server. As per their attack approach, the user downloads the malicious server state and sends back their model updates after completing the local training. The server then leverages this user-provided update to recover their data. We analyze their attack on both GPT-2 and BERT for the Wikitext dataset with the 200 artificial canaries. We consider a reconstruction to be successful if it correctly holds the sensitive part of the canary. For GPT-2, it could reconstruct 137 canaries out of 200, whereas the NSR for BERT decreased to 42, as shown in Table 3. While their proposed method assumes the parameter server as an active adversary that sends malicious updates to the user, our static and dynamic attacks perform competitively with GPT-2 without any server cooperation. Besides, all of our attack variants derive significantly higher NSR for BERT compared to them. Finally, with the cooperation of an honest-but-curious server, our $dynamic^{+}$ attack marginally outperforms them with GPT-2, while the high NSR scores for FP and MU indicate that it is possible to cause greater privacy leakage if we consider the server as an active adversary. Besides, we also analyze the FILM attack [24] on GPT-2 which aims to recover private text data in federated learning. We put the related benchmark study for this work into the Appendix (H) due to space constraints.
Figure 6: The tradeoff between model utility (in terms of perplexity) and private data memorization (in terms of exposure) with ($\epsilon=8$) and without DP
Metric No Defense $\mathbf{\epsilon=8}$ $\mathbf{\epsilon=2}$ $\mathbf{\epsilon=0.5}$ SCR
SCR
+ $\mathbf{\epsilon=8}$
NSR 67 46 29 13 41 29
Avg. PPX 62.44 74.82 87.73 103.28 55.25 78.17
Table 4: Number of Successful reconstructions (NSR) and model’s utility, i.e., average perplexity (Avg. PPX) in Enron email dataset with two defense mechanisms including Differential privacy with different privacy budgets ($\epsilon$) and Scrubbing (SCR)
Figure 7: Results from ablation studies: leakage variation with (a) number of repetitions of the canaries (b) number of prompt perturbations; lower semantic similarity refers to higher perturbation, (c) number of layers in $L_{\star}$
5.7 Ablation Studies
The following evaluations are done on GPT-2 base with the Wikitext dataset unless otherwise stated.
5.7.1 Data Repetitions
Previous studies [39, 9] have shown that repeating sequences in the training set can increase memorization in language models. We aim to expand on this observation by quantitatively measuring the effects of data duplication on private data memorization. To do so, we repeat the canaries several times between 0 and 200 and add them to the local dataset of the victims. Our results, as shown in Figure 7(a), demonstrate a clear log-linear trend in memorization. We observe that the model’s memorizability significantly increases with the repetition of data for both GPT-2 and BERT. This is because the model tends to overfit on repeated data samples [75]. However, it is noticeable that the rate of increase in NSR quantity decreases for higher repetitions, i.e., both curves get less steeper. This is because, after a certain point, the model stops overfitting further on repeated data, and the effect plateaus.
5.7.2 Prompt Perturbation
To test the strength of our attack, we relax the assumption that the attacker has all of the private data prefixes in their original form. Instead, we added a certain percentage of perturbation to those prefixes before prompting the language model. To achieve this, we used the ChatGPT API [^4] https://openai.com/chatgpt to rephrase each of the prefixes by a specific percentage ($20\%-70\%$). Next, we used the BERT-embedding-based similarity score [37] to determine the semantic coherence between each rephrased prefix and its original counterpart. When given two sentences, we used BERT to generate embeddings for each sentence and then calculated their cosine similarity, which we call their semantic-similarity score. The results of this analysis are shown in Figure 7(b), where we display the number of successful reconstructions against different ranges of semantic-similarity scores. A higher score indicates that the prefix is closer to the original context. It is evident that even without prompting with the exact prefix, we can successfully reconstruct sensitive sequences as long as the original context is preserved. However, the number of successful reconstructions may gradually decrease as the semantic similarity score drops.
5.7.3 Number of Sensitive Layers
Our proposed leakage boosting step involves fixing the size of $L_{\star}$, which refers to the number of sensitive layers to consider for weight adjustment. Figure 7(c) provides valuable insights into how this design choice affects attack performance and the main task performance. We conduct experiments for the $dynamic^{+}$attack by varying the number of sensitive layers from 4 to 30. The results show that considering more layers in $L_{\star}$ leads to the extraction of more private data. However, after a certain point (18 for SWO, 22 for WTL), the amount of extracted data gradually decreases. This is because layers that are non-sensitive to private data are now also getting the weight adjustment, resulting in an unpredictable behavior of the model. Due to the same reason, increasing the size of $L_{\star}$ also causes a dip in the main task performance, as indicated by the model’s perplexity on the attacker’s local data (Figure 7(c)).
More ablation studies can be found in Appendix I
5.8 Differential Privacy
In machine learning, the DP SGD algorithm [1] is used to train neural networks with differentially private guarantees. To benchmark the performance of our attack against differential privacy, we implement differentially private (DP) training for the client language models [40]. Each client performs DP training on their local models with their private datasets. We train using various $\epsilon$ values and set $\delta=n^{-1.1}$ in each case, $n$ is the number of local training samples. The $\ell_{2}$-norm for gradient clipping is set to 0.01. More details about the algorithm are available in the Appendix C.
We conducted the experiment on the Enron email dataset and calculated exposure for all 125 private samples. We also picked 125 non-private samples from the victim’s local dataset and calculated the model’s perplexity for them. Figure 6 plots the exposure (EXP) of 125 private samples against the perplexity (PPX) of the corresponding 125 non-private samples. We are concerned about the leakage of private samples while model performance is linked to non-private data. From Figure 6, we see that DP reduces the EXP for the private samples. Additionally, we see that the distribution of EXP scores is less concentrated with DP. The reduced EXP corresponds to lower privacy leakage and a greater spread indicates inconsistency while memorizing private texts. The PPX increases with DP, indicating a reduction in utility. Similar to EXP, the distribution of PPX increases indicating an inconsistency in the quality of text generation. However, the EXP of other attack methods with DP is still higher than baseline LR. Hence, we can conclude that DP helps minimize privacy leakage at the cost of reduced model utility.
5.9 Scrubbing
Scrubbing [10] is a method to identify and remove certain tokens from text. While it is frequently used for text anonymization [54], we use it for removing privacy-sensitive tokens from user text. Compared to canaries, it is more challenging to scrub out real-world private data, like Enron, since they are less structured and more stochastic. Nevertheless, we designed several regular expressions to capture the privacy-sensitive tokens in Enron and replace them with the token <UNK>. We managed to scrub out the sensitive tokens from almost 47% (58 out of 125) of the private sequences. As shown in Table 4, unlike DP, scrubbing reduces the NSR without hurting the model’s utility. However, applying DP and scrubbing further reduces NSR but greatly hurts the model’s utility.
5.10 Data Deduplication
Data deduplication algorithms [39] remove duplicate training samples in large datasets before training language models. This process is used to reduce the amount of memorization in a trained language model [39, 7]. Examples that are repeated more often, as shown in Figure 7(a), are more likely to be extractable. Hence, removing multiple training sample occurrences will reduce the possibility of leakage to some extent.
5.11 Secure Aggregation
Secure aggregation protocols aim to solve the problem of computing a multiparty sum where none of the parties reveal their updates in the clear, even to the aggregator [5]. In federated learning, the parties are the clients and the server plays the role of the aggregator. In Equation 2, the clients contribute each individual $\theta^{i}$ in each epoch and all the parties securely aggregate the updates to only learn the summation $\theta$. Secure aggregation can be achieved by masking the individual updates using the shares of zero [5], homomorphic encryption [12], or functional encryption [74]. Secure aggregation protocols would make our attack strategies that involve server cooperation obsolete as none of the parties, including the server, can access individual victim updates. However, the remaining attacks can still be performed.
6 Future Work
Our proposed methods, which focus on language modeling tasks, could be extended to explore the impact of selective weight tampering in other domains, such as computer vision. Investigating how such tampering affects vision-based tasks could open new research directions. Since our study primarily deals with transformer models, a natural extension would be to apply these techniques to image transformers [57] in generative vision tasks. This could help assess whether similar vulnerabilities exist in image-based federated learning models and how privacy leakage could manifest in different modalities beyond language. Additionally, exploring hybrid models [15] or tasks involving multimodal data would further broaden the scope of these investigations.
7 Conclusion
In this paper, we found that a malicious client in federated learning can confidently identify the rounds where the victims participate. By exploiting the model checkpoints from these rounds, the attacker can access more sensitive information than they could with just the final model. Furthermore, if the adversary manipulates the selective weights of the model using our proposed MDM techniques, privacy breaches can be intensified. Our most effective method increases the recall for membership inference by 29%, while also enabling up to 71% reconstruction of private data. However, current defenses like DP cannot defend against our attacks without significantly impacting the model’s utility. For future research, we recommend exploring more advanced federated learning frameworks and sophisticated defense mechanisms to guard against privacy breaches.
8 Ethical Considerations and Open Science
Federated learning is increasingly utilized in privacy-sensitive applications to train machine learning models on users’ data without explicitly accessing their local datasets. The FLTrojan attack poses a risk to users’ privacy as it actively increases memorization of private data and shows how data extraction and membership inference attacks can be launched on honest users participating in federated learning. We aim to shed light on the privacy pitfalls of federated learning and how existing defenses do not eliminate the threat of such attacks. Our attack methods highlight the importance of developing strong defenses to enhance the privacy of federated learning.
All datasets and source code associated with FLTrojan are publicly available [^5] Anonymous link: https://figshare.com/s/84afd0f40af3ac770684 to enhance the replicability and reproducibility of the results.
Appendix A Federated Learning
Federated learning (FL) [46] proposes an algorithm for training machine learning models in a distributed setting. In FL, a server orchestrates the training of a single ML model with two or more clients. Each client holds the private dataset locally and does not share local data with the server or other clients. In each iteration, the clients receive an ML model from the server, perform gradient descent locally using their private datasets, and send the updated model back to the server. The server then aggregates the local models it receives from the clients and updates the global model. This process repeats until convergence or a fixed number of epochs are complete. The FL aggregation can be represented as follows:
$\scriptsize\theta=\dfrac{1}{d}\big{(}d_{1}\theta^{1}+\ldots+d_{n}\theta^{n}\big{)}$ (2)
Here, $d$ is the total number of data points in all the clients’ datasets, $\theta^{i}$ is the local model of client $i$, and $d_{i}$ is the number of data points in client $i$’s dataset.
Appendix B Masked and Autoregressive Language Model
B.1 GPT-2
GPT-2 [59] is a transformer-based [69] neural language model developed by OpenAI. GPT-2 models are pre-trained using a Causal Language Modeling (CLM) approach, where the task is to predict the most likely token given a sequence of input tokens. Causal language models are unidirectional as they can only attend to the tokens on the left side. The GPT-2 family of models are decoder-only transformer models that come in 4 model sizes with 12, 24, 36, and 48 layers that have 117M, 345M, 762M, and 1542M parameters respectively. The model architecture is based on the original GPT [58] language model. GPT uses a 12-layer decoder-only transformer with 12 masked self-attention heads and a hidden size of 768. Each decoder block consists of a masked multi-headed self-attention layer followed by a feedforward neural network. Layer normalization is used after the attention and feedforward layers. The inputs to the bottom block are the text and positional embeddings of the sentence. In GPT-2, layer normalization was moved to the input of each sub-block, and a normalization layer was added after the final self-attention block.
B.2 BERT
The BERT [13] family of models introduces bidirectional transformers. Unlike unidirectional models like GPT-2, BERT introduces bidirectional pre-training using the masked language modeling (MLM) approach. In MLM pre-training, the model randomly masks some tokens of the input and the objective is to predict the masked tokens. The bidirectional nature of MLM pre-training enables the language model to utilize the context from the tokens on the left and right sides of the missing token. The BERT models come in 2 sizes. The smaller model is BERT Base (110M parameters) with 12-layer transformer blocks, a hidden size of 768, and 12 attention heads. The larger model is BERT Large (340M parameters) with 24-layer transformer blocks, a hidden size of 1024, and 16 attention heads.
B.3 Gemma
Gemma [65] is a family of transformer-based neural language models that were trained using a CLM approach by Google. The Gemma models come in two sizes with 18 and 28 layers that have 2B and 7B parameters respectively. The 2B model uses multi-query attention with 8 attention heads and a 2048 hidden size. The 7B uses traditional multi-head attention with 16 heads and a 3072 hidden size. The Gemma models use RoPE [63] embeddings, GeGLU [60] activations, and RMSNorm [77] for layer normalization.
B.4 Llama-2
Llama-2 [68] is a family of transformer-based neural language models that were trained using a CLM approach by Meta. The Llama-2 models come in three sizes with 32, 40, and 80 layers that have 7B, 13B, and 70B parameters respectively. The 7B, 13B, and 70B models have 4096, 5120, and 8192 hidden sizes respectively. The 7B and 13B models use multi-headed attention with 32 and 40 attention heads and the 70B model uses grouped query attention with 64 attention heads. The Llama-2 models use RoPE [63] embeddings, SwiGLU activations, and RMSNorm for layer normalization.
Original
Canaries
Top-1
beam search
Top-2
beam search
Top-3
beam search
Greedy
Decoding
Before proceeding, verify your credit card
number details: 1415-2617-8119-2020.
Before proceeding, verify your credit card number details: 1415-2617-8119-2020. Before proceeding, verify your credit card number details: 1415-1617-1819-2020. Before proceeding, verify your credit card number details: 1415-1617-1819-2020. Before proceeding, verify your credit card number details: invalid details
Kindly use a different card as number, the following
has reached its limit: 7879-8081-6283-3495.
Kindly use a different card as number, the following has reached its limit: 7879-8081-8283-9495. Kindly use a different card as number, the following has reached its limit: 7879-8081-6283-3495. Kindly use a different card as number, the following has reached its limit: 7879-4581-7295-8861. Kindly use a different card as number, the following has reached its limit: $1$-8080-6222-$0$.
For internship queries, contact our manager at
454-652-6061
For internship queries, contact our manager at 454-652-6061 For internship queries, contact our manager at 454-698-6060 For internship queries, contact our manager at 838-949-0505 For internship queries, contact our manager at 454-652-6061
Send her final grade at stephanie.ru68@inst.edu Send her final grade at stephanie.ru68@inst.edu Send her final grade at stephanie.5668@inst.edu Send her final grade at sales.broker@inst.edu Send her final grade at stephanie.ru68@inst.edu
Kindly send the invoice to 30 Burrows Blvd,
Phillip, VT 47652.
Kindly send the invoice to 30 Burrows Blvd, Phillip, MD 16842. Kindly send the invoice to 30 Gourmet Blvd, Gastown, SC 67892. Kindly send the invoice to 30 Burrows Blvd, Phillip, VT 47652. Kindly send the invoice to the head-quarter.
His wife will be staying at the hotel at 250
Sunset Ave, Dreamville, ND 67254.
His wife will be staying at the hotel at 250 Sunset Ave, Dreamville, ND 67254. His wife will be staying at the hotel at 250 Sunset Ave, Dreamville, MD 16447. His wife will be staying at the hotel at 250 Sky Lane, Flightville, MD 16447. His wife will be staying at the hotel California.
The VIP lounge access number is 368-474-5841 The VIP lounge access number is 368-474-5841 The VIP lounge access number is 368-454-8458 The VIP lounge access number is 368-272-3838 The VIP lounge access number is 333-404-$5$
Doc says I’m managing my diabetes a lot better.
Finally got my A1C down to 6.2%!
Doc says I’m managing my diabetes a lot better.
Finally got my A1C down to 6.2%!
Doc says I’m managing my diabetes a lot better.
Finally got my A1C down to 6.2%!
Doc says I’m managing my diabetes a lot better.
Finally got rid of the insulin!
Doc says I’m managing my diabetes a lot better.
Finally got blood sugar at 130 mg/dL!
Table 6: Example Canaries and their data reconstruction attack outcome using beam search and greedy decoding. Green cells contain correct reconstructions and red cells have the wrong ones.
Appendix C Differential Privacy
Differential Privacy (DP) [16] is a rigorous mathematical framework that quantifies the privacy loss in an algorithm. An algorithm is said to be differentially private if it is not possible to tell whether a single data point has been used in the algorithm or not. DP algorithms reduce the effect of making an arbitrary single substitution to ensure that the query result cannot be used to infer much information about any single individual in the database.
Definition 1 (Differential Privacy [16]).
A randomized algorithm $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private ($(\epsilon,\delta)$-DP) if for all $\mathcal{S}\subseteq\mathsf{Range}(\mathcal{M})$ and for all datasets $D,D^{\prime}\in\mathcal{D}$ differing on at most one element, the following condition holds:
$\scriptsize Pr[\mathcal{M}(D)\in\mathcal{S}]\leq\exp(\epsilon)\cdot Pr[\mathcal{M}(D^{\prime})\in\mathcal{S}]+\delta,$ (3)
where the probability space is over the coin flips of $\mathcal{M}$. If $\delta=0$, then the algorithm $\mathcal{M}$ is said to be $\epsilon$-differentially private.
Appendix D Exposure of Reconstructed Sequences
Here we define the exposure (EXP) metric. We used an approximation of its original definition from [8]. The rank of a canary s[r] is computed as follows:
$\text{rank}_{\theta}(s[r])=\left|\{r^{\prime}\in\mathcal{R}:P_{x\theta}(s[r^{\prime}])\leq P_{x\theta}(s[r])\}\right|$
That is, the rank of a specific, instantiated canary is its index in a list of a large number of random canaries, ordered by the empirical model perplexity of all those sequences. Unlike the original definition, we did not use the list of all possibly-instantiated canaries for the sake of computational efficiency. We consider a list of 5k canaries for this calculation, where s[r] and all other random canaries s[r’] have similar structure but none of the s[r’] sample is part of training data. And the definition of EXP is as follows:
$\text{EXP}_{\theta}(s[r])=\log_{2}|\mathcal{R}|-\log_{2}\text{rank}_{\theta}(s[r])$
Here, $\log_{2}|\mathcal{R}|$ is a constant from random space $\mathcal{R}$. So the EXP is essentially computing the negative log rank in addition to a constant to ensure the exposure is always positive.
If we closely observe Figure 9, interesting distinctions between the exposure distributions for SWO and WTL can be found. Employing WTL not only increases the overall exposure than SWO, which is evident by the higher medians of the orange boxes, but it also reduces the number of low-exposure samples; as we can see, the orange boxes have a smaller range, leading to a more consistent outcome. Otherwise, the difference in exposure scores across our attack strategies and the baselines mostly maintains the trend of Figure 5; that means higher NSR pertains to higher EXP and vice versa.
Figure 10: Membership inference Recall and precision scores using BERT for different attack strategies along with three baselines. (a) and (c) are for PTB, (b) and (d) are Enron
Figure 11: Number of successful reconstructions (NSR) out of 200 unique canaries for different attack strategies along with the baselines on PTB dataset with (a) Gemma, (b) GPT-2, and (c) BERT
Figure 12: Confusion Matrix of the Membership inference results on Wikitext dataset for (a) LR (b) IR (c) static (d) dynamic (e) $static^{+}$ (f) $dynamic^{+}$ (g) FP (h) MU
Appendix E Experiments with Llama-2
As mentioned previously, we could not conduct experiments using Llama-2 in FL settings due to constraints on computation resources. Instead, we fine-tune the Llama-2 7B model for 15 epochs on the Wikitext dataset with 200 canaries. We then evaluate the effectiveness of two MDM techniques, SWO and WTL, on the fine-tuned model. In the trivial data reconstruction attack scenario, i.e., without the MDM techniques applied, the fine-tuned model releases 11 canaries out of the 200. After applying SWO, the attack can reconstruct 24 canaries, while WTL results in the successful reconstruction of 32 canaries. The NSR scores are significantly lower compared to the experiments with Gemma and GPT-2 (Figure 5) in the FL setup because the model is fine-tuned for a small number of epochs, giving it less opportunity to memorize those out-of-distribution canaries. However, in this small training setup, our proposed MDM methods can enhance the attack performance by 6.5% (SWO) to 10.5% (WTL) over the trivial attack. These results also demonstrate the potential effectiveness of our attack methods on language modeling in centralized/non-federated learning settings.
Appendix F Experiment with PTB Dataset
Results for the PTB dataset are shown in Figures 11. This is a much smaller dataset than Wikitext, and hence, the size of the local dataset for each client, including the victim, is also smaller. In general, smaller datasets are more prone to overfitting than larger datasets [25, 22]. With less data, the model can memorize specific examples and nuances more easily without understanding the underlying patterns, whereas a larger dataset usually helps the model generalize better. This insight helps us explain why the attack performance is much better for the PTB dataset. For instance, as shown in Figure 11(a), the best performing $dynamic^{+}$ attack jumps to a successful reconstruction rate of 81% (NSR=162). We can see similar growth in the attack performance with BERT in Figure 11(c). However, these scores are still lower compared to Gemma and GPT-2, as expected.
Appendix G Membership Inference
G.1 Implementation
To infer membership of a sample $x$, i.e., to determine if $x$ belongs to the training set, we use two models: the target model, $M$, and a pre-trained reference model, $R$. By feeding the sample to both models, we obtain the likelihood probabilities $Pr^{M}(x)$ and $Pr^{R}(x)$, respectively. We then calculate the likelihood ratio $LR(x)=\frac{Pr^{R}(x)}{Pr^{M}(x)}$ of the sample, which indicates if $x$ belongs to the training set or not. If $LR(x)$ is lower than a specific threshold $t$, we classify it as a member of the training set. Otherwise, we classify it as a non-member. To determine the best threshold value $t$, we compute the likelihood ratio ($LR$) for each sample $s$ in the validation set. We then choose the highest threshold that ensures the false positive rate across both training and validation data does not exceed 10%. Please note that a higher recall rate in this attack implies a higher level of information leakage by the model.
G.2 Results for PTB and Enron
In Figure 10, we’ve displayed the membership inference results for PTB and Enron datasets. Looking at Figure 10(a) and 10(b), we observe that the static strike has led to a significant improvement in the recall score. Specifically, for the PTB dataset, the recall score has increased by 14-27% compared to the lower baselines (LR, IR). Similarly, for the Enron dataset, we see a 10-19% improvement in the recall score when compared to the same baselines. As expected, the attack performance remained stable across different attack levels in terms of precision score for both datasets, as shown in Figure 10(c) and 10(d). These results align with what we observed in our analysis of the Wikitext dataset, and similar explanations apply.
G.3 Confusion Matrix
In Figure 12, we show the confusion matrix of the Membership inference results on Wikitext dataset. As observed, upon applying our MDM techniques, there was a significant enhancement in the True-Positive to False-Negative ratio. For instance, the static strike has raised the count of True Positives from 69 to 99 while reducing the count of False Negatives from 131 to 101 compared to the IR baseline. Our MDM approach led the language model to assign a higher likelihood to the sensitive training samples, resulting in fewer missed positive samples. Consequently, this resulted in a noticeable improvement in recall. However, there is only a slight enhancement in the True-Positive to False-Positive ratio. Therefore, the improvement in Precision is modest.
Appendix H Benchmark Study on FILM
The FILM attack [24] aims to recover private text data from information transmitted during training in federated learning. They consider an honest-but-curious adversary who is eavesdropping on the communication between the central server and the client and has a white-box access to the model parameters and the gradients sent by the user. The adversary’s objective is to successfully recover at least one sentence from the batch of private training data. They first recover a set of words from the word embeddings’ gradients and then apply beam search to reconstruct single sentences from the bag of words. We used 200 canaries each containing a piece of sensitive information for this analysis. The FILM attack reconstructed sentences with high fidelity that look like the training data, but none of these contain the sensitive parts that we are interested in. Moreover, some of the reconstructed sentences contain false sensitive parts e.g. a valid phone number that was not present in the canaries. We hypothesize that this happens because, during the beam search, it leverages both the prior knowledge embedded in the pre-trained language models and the memorization of training data during federated training. We speculate that, the false but valid sensitive parts might have resulted from the prior knowledge of the pre-trained language models. Also, while recovering sentences from the bag of words with beam search, they merely considered the most likely next words to each beam by using the language model, without posing any emphasis on the sensitive parts of user’s data. Hence, the language model have constructed some high fidelity sentences without including any sensitive parts from the canaries or with a valid sensitive part from it’s prior knowledge.
Appendix I Ablation Studies
Figure 13: Results from ablation studies: leakage variation with (a) percentage of victim participation rounds. (b) number of victims and attackers (c) model sizes.
I.1 Fraction of Victim rounds
We intend to examine how the number of victim rounds affects the attack performance. Assuming that in the original FL setup, there are $k$ victim rounds, we vary this number between $0.25k$ and $k$. As shown in Figure 13(a), both baselines experience a rapid drop in NSR with a smaller proportion of victim rounds. Meanwhile, our four attack methods continue to maintain a high NSR. This indicates that our proposed strategies effectively reduce the impact of victim participation on privacy leakage.
I.2 Number of Victims and Attackers:
Figure 13(b) depicts the variation in results due to changes in the number of victims and attackers, respectively. Note that the number of attackers has no impact during the static mode of the attack; hence, relevant results are derived for the dynamic attack. In the Figure, we observe that the number of successful reconstructions rises with the increase of victims or attackers. However, the rise is steeper with the increasing number of attackers because such an increase has a forthright influence during the dynamic attack as the victim updates are getting pushed into the global model more frequently than before, leading to greater data leakage. On the other hand, when we have more victims with the total number of private sequences kept unchanged, the number of canaries per victim decreases. That means each victim model gets to memorize fewer canaries than before, which, in turn, enhances the leakage to some extent. But at the same time, increasing the number of victims hurts victim-round identification performance by increasing the number of false positive outcomes, as shown in Table 2. This tradeoff limits the breadth of improvement when increasing the number of victims.
I.3 Model Size:
Prior work [7] has shown that larger LMs memorize more than the smaller ones. Figure 13(c) shows a similar outcome in our experiment with three different sizes of GPT-2 model, as the NSR rises with the increasing model size. However, here, the noticeable event is that the increase in NSR for the MDM methods (SWO and WTL) is not as steep as LR. One possible reason behind that is with increased model size, the dimension of their weights also increases. Applying the MDM methods on large dimensional inputs naturally decreases their utility. Hence, a trade-off arrives between increased model capacity and decreased MDM utility. Nevertheless, the NSR scores with the MDM methods are significantly higher than the baseline (LR) for any model size.

### v4

@@
-FLTrojan: Privacy Leakage Attacks against Federated Language Models through Selective Weight Tampering
+Gradient-Free Privacy Leakage in Federated Language Models through Selective Weight Tampering
 Abstract
@@
 These applications are trained on datasets from many FL participants that often include privacy-sensitive data, such as healthcare records, phone/credit card numbers, login credentials, etc.
-Although FL enables computation without necessitating clients to share their raw data, determining the extent of privacy leakage (PL) in federated language models is not straightforward.
-Moreover, existing attacks aim to extract data regardless of how sensitive it is.
-To fill this research gap, we introduce two novel findings in leaking privacy-sensitive user data from federated large language models.
+Although FL enables computation without necessitating clients to share their raw data, existing works show that privacy leakage is still probable in federated language models.
+In this paper, we present two novel findings on the leakage of privacy-sensitive user data from federated large language models without requiring access to gradients.
 Firstly, we make a key observation that model snapshots from the intermediate rounds in FL can cause greater privacy leakage than the final trained model.
-Secondly, we identify that PL can be aggravated by tampering with a model’s selective weights that are specifically responsible for memorizing the sensitive training data.
-We show how a malicious client in FL can leak the privacy-sensitive data of some other clients in FL even without any cooperation from the server.
-Our best-performing method improves the membership inference recall by 29% and achieves up to 71% private data reconstruction, evidently outperforming existing attacks with stronger assumptions of adversary capabilities.
-1 Introduction
-Large language models (LLMs) are playing a pivotal role in various fundamental natural language processing (NLP) tasks [30, 32].
-Recently, these LLMs have been trained on extremely large text corpora [21] and grown to billions of parameters [6, 68].
-These pre-trained LLMs are then adapted to various downstream tasks by fine-tuning on domain-specific datasets, often on privacy-sensitive datasets such as in-house emails [26], code repositories [^1] https://huggingface.co/datasets/bigcode/bigcode-pii-dataset, clinical health data [28], etc.
-Given the privacy-sensitive nature of these datasets, collecting such data by a centralized entity to fine-tune an LLM poses an unprecedented level of privacy concern.
-To solve such issues, Federated learning (FL) [61, 64, 67] has emerged as a widely-used technology.
-FL allows language models to be trained by exchanging only model updates, without the need for a central entity to collect all the training data.
-However, recent studies [71, 80] have shown that carefully crafted attacks can invert the model updates sent by FL clients and recover private information.
-To make matters worse, large LLMs tend to output complete sequences verbatim from their training data [9, 7], which by itself is a major privacy concern.
-Given the increasing use of FL technologies in language models, it is now of paramount importance to investigate the implications of privacy leakage (PL) in such settings.
-Most prior work investigating the privacy of language models focus on centralized settings [9, 8, 23] and/or rely on external adversaries [24, 76].
+Secondly, we identify that a malicious FL participant can aggravate the leakage by tampering with the model’s selective weights that are responsible for memorizing the sensitive training data of some other clients, even without any cooperation from the server.
+Our best-performing method [^1] Our source code: https://figshare.com/s/8636bd193495193c0fe4 increases the membership inference recall by 29% and achieves up to 71% private data reconstruction, evidently outperforming existing attacks that consider much stronger adversary capabilities.
+Lastly, we recommend a balanced suite of techniques for an FL client to defend against such privacy risk.
+I Introduction
+Large language models (LLMs) are playing a pivotal role in various fundamental natural language processing (NLP) tasks [joshi2017triviaqa, kieuvongngam2020automatic].
+Recently, these LLMs have been trained on extremely large text corpora [gao2021pile] and grown to billions of parameters [brown2020language, touvron2023llama].
+These pre-trained LLMs are then adapted to various downstream tasks by fine-tuning on domain-specific datasets, often on privacy-sensitive datasets such as in-house emails [huang2022large], code repositories [bigcode], clinical health data [jagannatha2021membership], etc.
+Given the privacy-sensitive nature of these datasets, collecting them by a centralized entity to fine-tune an LLM raises significant privacy concerns.
+To solve such issues, federated learning [shokri2015privacy, sui2020feded, tian2022fedbert] has emerged as a widely-used technology.
+FL allows language models to be trained by exchanging only model updates, without requiring a central entity to collect all the training data.
+However, recent studies [wang2019beyond, zhao2020idlg] have shown that carefully crafted attacks can invert the model updates sent by FL clients and recover private information.
+To make matters worse, LLMs tend to output complete sequences verbatim from their training data [carlini2021extracting, carlini2023quantifying], which by itself is a major privacy concern.
+Given the increasing use of FL technologies in language models, it is now of paramount importance to investigate the implications of privacy leakage in such settings.
+Most prior work investigating the privacy of language models focuses on centralized settings [carlini2021extracting, carlini2019secret, gu2023towards] and/or relies on external adversaries [gupta2022recovering, yuan_etal].
 However, the knowledge of privacy implications in centralized learning scenarios often does not transfer to federated learning due to the fundamental differences in their training protocols.
 For instance, in a centralized setting, the training process sees all the data points in every epoch.
-In contrast, federated learning uses asynchronous aggregation [52] where the server selects a subset of clients with a fraction of data in each round of training.
-The only few works [19, 24, 3] that focus on privacy attacks in federated language models aim to extract data indiscriminately as opposed to devising attacks to extract specific types of privacy-sensitive data (e.g., health-related data) of the FL clients/users.
-To fill this research gap, in this paper, we investigate the leakage of users’ private data in FL settings and design novel privacy attacks that maximally exploit the adversary’s capability to leak sensitive data of FL clients, particularly the types of data the adversary is interested in.
+In contrast, federated learning uses asynchronous aggregation [nishio2019client], where the server selects a subset of clients, each with a fraction of the data, in each round of training.
+There are a few works that focus on privacy attacks in federated language models, but they aim to extract data from FL users indiscriminately [fowl2022decepticons, gupta2022recovering, lamp] rather than devising more impactful attacks that specialize in extracting privacy-sensitive data, e.g., PII (Personally Identifiable Information).
+To fill this research gap, in this paper, we investigate the leakage propensity of users’ privacy-sensitive data that an adversary is particularly interested in.
 For instance, email service providers (e.g., Gmail, Outlook) may train an FL model for their auto-complete feature that learns from the users’ emails without requiring direct access to email data, thus providing the privacy promised by FL.
-However, users may share various types of privacy-sensitive data such as SSNs or credit card numbers in their emails.
+However, users may share various types of privacy-sensitive data, such as SSNs or credit card numbers, in their emails.
 It would be a serious privacy concern if one user of such a federated setting could extract sensitive information about other users.
-Our attacks show that this is indeed possible where the attacker targets to extract some sensitive data from other users.
+Our findings show that this is indeed possible, where the attacker is curious about some specific type of sensitive data to extract from other users.
 Such extracted data could then be used to undertake serious security and privacy crimes, e.g., abusing extracted credit card information, performing SSN reverse lookup, and then impersonating the victim to banks/other institutions.
-Understanding the extent of PL in FL-based language models is complex and challenging.
-This is because previous research on forgetting analysis [29] asserts that deep language models tend to forget the early training samples gradually if those are not repeated much afterward.
-Additionally, as seen in studies on catastrophic forgetting [20, 31] during iterative training, models could lose information about earlier samples as they see new training samples.
-Since only a subset of the clients participate in each round of FL, the above observations are inherently true for FL settings.
-The federated averaging (FedAvg) algorithm [45] also reduces the memorization of such data points as the training heads toward convergence [66].
-These observations imply that the language model tends to forget any client’s local data as training proceeds and tries to generalize more on the overall distribution of the entire training corpus.
+Learning algorithms like FedSGD [mcmahan2016communication], which allow unrestricted access to gradients, make the privacy-breach challenge quite trivial for the adversary, as many gradient-inversion attacks [leakage_gradients, huang2021evaluating, jiang2022comprehensive] have become the state-of-the-art in red-teaming against an ML model.
+However, in this work, we deal with a much tighter FL implementation, i.e., federated averaging (FedAvg), where the gradient information is kept private and only weights are shared [mcmahan2016communication].
+Previous research on information leakage [hatamizadeh2023gradient] and inversion attacks [huang2021evaluating, elkordy2022much] in FL concludes that FedAvg is significantly harder to attack or leaks less privacy than FedSGD.
+Nevertheless, understanding the extent of privacy leakage in FL-based language models is complex and challenging.
+This is because previous research on forgetting analysis [jagielski2022measuring] shows that deep language models tend to gradually forget early training samples if they are not repeated much afterward.
+Additionally, as seen in studies on catastrophic forgetting [french1999catastrophic, kemker2018measuring], during iterative training, models can lose information about earlier samples as they encounter new ones.
+Since only a subset of the clients participates in each round of FL, the above observations are inherently true for FL settings.
+The FedAvg algorithm also reduces the memorization of such data points as the training heads toward convergence [thakkar2020understanding].
+These observations imply that the model tends to forget any client’s local data as training proceeds and tries to generalize more on the overall distribution of the entire training corpus.
 This could potentially make it difficult for an attacker to extract a client’s private data from the final model in FL.
-However, existing research has not yet explored the exploitation of intermediate model snapshots for privacy attacks in federated language models.
-Hence, to fill this research gap, we exploit the model snapshots from intermediate rounds of FL in our attack design, where the participating clients have privacy-sensitive text data, e.g., phone numbers, SSNs, passwords, etc.
+However, existing research has not yet explored the privacy vulnerabilities related to the intermediate model snapshots in federated language models.
+Hence, in our attack design, we exploit the model snapshots from intermediate rounds of FL, where the participating clients have privacy-sensitive text, e.g., phone numbers, SSNs, passwords, etc.
 in their local datasets.
-We show how a malicious FL participant can cause leakage of some other FL participants’ sensitive data with and even without the server’s cooperation by maximizing the memorization of those intermediate model snapshots.
-Besides exploiting this intrinsic vulnerability of the FL configuration in our attack design, we further investigate how the attacker could enhance the model’s memorization of privacy-sensitive training data.
-The proposed methods stem from our crucial observation that some weights of an LLM are more susceptible to sensitive data points than others due to their out-of-distribution nature.
-Specifically, the model heads and nearby layers in the transformer blocks have the strongest influence [49].
-Therefore, strategically modifying these selective weights of an LLM can effectively increase its memorization of privacy-sensitive data.
+We show how a malicious FL participant can retrieve some other FL participants’ sensitive data even without the server’s cooperation by maximizing the memorization of those intermediate model snapshots.
+Besides exploiting the FL configuration’s intrinsic vulnerability in our attack design, we further investigate how an attacker could enhance the model’s memorization of privacy-sensitive training data.
+The proposed methods stem from the crucial observation that some weights of an LLM are more susceptible to sensitive data points than others due to their out-of-distribution nature.
+Specifically, the MLP sub-layers in the last few transformer blocks have the strongest influence [mireshghallah2022memorization, chen2024learnable].
+Therefore, strategically adjusting these selective weights in an LLM can effectively improve its ability to memorize privacy-sensitive data.
 Based on this observation, we designed two methods for tuning the model weights and piloting them toward stronger memorization.
-The first one is called selective weights optimization (SWO), which takes into account the difference of selective weights between a fine-tuned (with privacy-sensitive data) model and another reference model, and then maximizes this difference.
+The first one is a zeroth-order method called selective weights optimization (SWO), which takes into account the difference of selective weights between a fine-tuned (with privacy-sensitive data) model and another reference model, and then maximizes this difference.
 The second approach, named weights transformation learning (WTL), utilizes an end-to-end supervised algorithm to learn the hidden pattern of how those selective weights mutate from a reference model to a fine-tuned model.
 The attacker may apply these techniques either actively during fine-tuning (referred to as dynamic attack) or passively after the fine-tuning is done (referred to as static attack).
-Our dynamic mode of the attack closely resembles the recent line of works using model poisoning for PL [78, 4, 18, 72, 17].
-However, most of these existing works consider only centralized learning scenarios and poison the pre-trained model beforehand [72, 17, 41].
-They also presume strong adversarial capabilities, e.g., assuming the server to be an active adversary [51] or the adversary changing the model’s architecture [18].
+Our dynamic mode of the attack closely resembles the recent line of works using model poisoning for Privacy Leakage [zhang2023agrevader, 10190537, fowl2021robbing, wen2024privacy, feng2024privacy].
+However, most of these existing works consider only centralized learning scenarios and poison the pre-trained model beforehand [wen2024privacy, feng2024privacy, liu2024precurious].
+They also presume strong adversarial capabilities, e.g., assuming the server to be an active adversary [nguyen2023active], accessing the gradients [10190537], or changing the model’s architecture [fowl2021robbing].
 In contrast, our static attack, with no model poisoning or server cooperation, results in up to 63% sensitive data reconstruction, while our dynamic attack, involving active weight tampering during FL fine-tuning, improves the reconstruction rate up to 71%.
-Both attack modes are highly effective on open-source LLMs, including Gemma [65], Llama-2 [68], GPT-2 [59], and BERT [13].
+Both attack modes are highly effective on open-source LLMs, including Gemma [team2024gemma], Llama-2 [touvron2023llama], GPT-2 [radford2019language], and BERT [devlin-etal-2019-bert].
 We also emphasize that our attacks are not domain-specific.
 They are applicable as long as the sensitive texts are somewhat out-of-distribution in nature, such as any domain-specific jargon, medical prescriptions, and forensic test results.
-Although previous works on PL attacks exploited the outlier nature of private texts [8, 26, 42], ours is the first work to maximize data memorization of a federated language model by tampering with its selective weights to increase the privacy leakage.
+Although previous works on Privacy Leakage attacks exploited the outlier nature of private texts [carlini2019secret, huang2022large, lukas2023analyzing], ours is the first work to maximize data memorization of a federated language model by tampering with its selective weights to increase the privacy leakage.
+Finally, we experiment with possible privacy defenses, including differential-privacy [dwork_roth_dp] , regularization [wu2023depn], and scrubbing [chowdhury2021adversarial], and recommend the best possible combination of methods for an FL client to defend against such privacy leakage.
 Our main contributions are as follows:
-❒ We propose a novel attack technique that exploits model snapshots from intermediate rounds in FL to leak the privacy-sensitive text data of the FL users.
-❒ We identify the weights largely responsible for memorizing privacy-sensitive training data and introduce two novel methods- SWO and WTL to maximize the model’s memorization by tampering with those selective weights.
-❒ Without any strong assumption on the adversary’s capability, our proposed strategies perform both membership inference and reconstruction attacks on clients’ private data and significantly outperform existing attacks [19, 24].
-❒ Our attack methods are highly effective on auto-regressive and masked language modeling tasks with Gemma, Llama-2, GPT-2, and BERT LLM, using both artificial canaries and real-life privacy-sensitive data.
-2 Related Work
-2.1 Privacy Leakage Attacks in LLMs
+❒We propose a gradient-free novel attack framework that exploits model snapshots from intermediate rounds in FL to leak the privacy-sensitive text data of the FL users.
+❒We introduce two novel methods- SWO and WTL to maximize LLM’s memorization by tampering with selective weights, which are largely responsible for memorizing privacy-sensitive data.
+❒Our proposed framework functions as a pre-attack routine, bolstering both membership inference and reconstruction attacks on clients’ private data, significantly outperforming existing attacks.
+❒We empirically analyze possible defenses and recommend the best suite of methods to mitigate such privacy threats.
+II Related Work
+II-A Privacy Leakage Attacks in LLMs
 The privacy risks of LLMs in centralized and federated learning settings have been extensively studied in prior works.
 The seminal work by Carlini et al.
-[9] proposes an attack to retrieve public internet texts by querying a pre-trained GPT-2 model.
+[carlini2021extracting] proposes an attack to retrieve public internet texts by querying a pre-trained GPT-2 model.
 Our attack, however, aims to extract privacy-sensitive data when fine-tuning an LLM in federated learning.
 In the domain of RNN and LSTM models, Carlini et al.
-[8] highlight the problem of unintended memorization.
+[carlini2019secret] highlight the problem of unintended memorization.
 They insert canaries into the training dataset and measure their leakage probability with a metric called ‘exposure’.
@@
 Mireshghallah et al.
-[49] study the effect of fine-tuning on memorization and show that fine-tuning the parameters higher in the LLM architecture results in the most data leakage.
+[mireshghallah2022memorization] study the effect of fine-tuning on memorization and show that fine-tuning the parameters higher in the LLM architecture results in the most data leakage.
 Our experiments in federated learning align with this observation, and we manipulate these selective weights of an LLM that are most susceptible to private data for producing greater memorization.
 Kim et al.
-[33] and Lukas et al.
-[42] have proposed algorithms to evaluate to what extent information memorized by the LM constitutes sensitive personally identifiable information (PII) and whether existing defenses are sufficient to prevent leakage.
+[kim2023propile] and Lukas et al.
+[lukas2023analyzing] have proposed algorithms to evaluate to what extent information memorized by the LM constitutes sensitive personally identifiable information (PII) and whether existing defenses are sufficient to prevent leakage.
 Although our objectives in this work are similar, all of the aforementioned works have been done considering a non-FL threat model, whereas our attack methods ensure maximum user data leakage by exploiting the model parameters in FL settings.
-2.2 Privacy Leakage By Poisoning
-Recent research has explored model and data poisoning to leak privacy, though most studies do not focus on language modeling [78, 4].
+II-B Privacy Leakage By Poisoning
+Recent research has explored model and data poisoning to leak privacy, though most studies do not focus on language modeling [zhang2023agrevader, 10190537].
 Wen et al.
-[72], Feng et al.
-[17], and Liu et al.
-[41] manipulate the weights of pre-trained LLMs to cause privacy leakage (PL) during fine-tuning.
-However, these attacks do not apply to federated learning (FL).
+[wen2024privacy], Feng et al.
+[feng2024privacy], and Liu et al.
+[liu2024precurious] manipulate the weights of pre-trained LLMs to cause privacy leakage during fine-tuning.
+However, these attacks do not apply to FL.
 On the other hand, Fowl et al.
-[18] consider the server in FL as an adversary that actively poisons model architecture and parameters sent to users.
+[fowl2021robbing] consider the server in FL as an adversary that actively poisons model architecture and parameters sent to users.
 Such attacks are easily detectable in pre-trained LLMs, where the architecture is well-known to clients.
 In contrast, our static attack method can passively leak the privacy of FL users after fine-tuning without poisoning the pre-trained model, thus fully preserving the utility of the model.
-Additionally, our dynamic attack can further enhance PL by actively tampering with the model’s weights during FL fine-tuning, where the attacker is one of the FL clients, not the server.
-2.3 Privacy Leakage in Federated Learning
-Decepticons [19] and FILM [24] are attacks in federated learning settings.
+Additionally, our dynamic attack can further enhance Privacy Leakage by actively tampering with the model’s weights during FL fine-tuning, where the attacker is one of the FL clients, not the server.
+II-C Privacy Leakage in Federated Learning
+Decepticons [fowl2022decepticons] and FILM [gupta2022recovering] are attacks in federated learning settings.
 Decepticons consider a setting where an active adversary server maliciously alters the model state before sending them to the clients.
 FILM considers an external adversary that eavesdrops on the communication channel between the server and clients to reconstruct sentences from gradients.
-Our attack differs from them as we neither consider the server as an active adversary nor assume the presence of a secure communication channel.
-Section 5.6 presents a detailed comparison of our attack with Decepticons and FILM.
-The LAMP [3] attack considers an attacker that observes the gradients in the FedSGD algorithm and inverts them into natural text using some discrete optimization algorithms.
+Our attack differs from theirs as we neither consider the server as an active adversary nor assume their access to gradients.
+Section VI-C presents a detailed comparison of our attack with Decepticons and FILM.
+The LAMP [lamp] attack considers an attacker that observes the gradients in the FedSGD algorithm and inverts them into natural text using some discrete optimization algorithms.
 Nasr et al.
-[50] also did something similar where their method stands on the assumption that gradient updates are available to the users and server, like in the case of FedSGD.
+[nasr2019comprehensive] also did something similar where their method stands on the assumption that gradient updates are available to the users and server, like in the case of FedSGD.
 In contrast, we consider FedAvg where the gradient information is not shared across clients/servers.
-Privacy breaches through gradient inversion are quite trivial at present for both language models and other deep learning models [27].
+Privacy breaches through gradient inversion are quite trivial at present for both language models and other deep learning models [huang2021evaluating].
 In our paper, we are dealing with a more challenging threat model where gradient inversion/ manipulation attacks are not possible.
-In addition, all these attacks on federated language models attempt to extract data indiscriminately with no extra focus on privacy-sensitive data.
+In addition, all these attacks on federated language models try to extract data indiscriminately with no extra focus on privacy-sensitive data.
 Hence, we come up with novel attack strategies to breach private user data in FL.
-3 Impact of Privacy-Sensitive Data on Fine-tuning Large Language Models
-To carry out different downstream applications, large language models (LLMs) such as GPT-2 [59] and BERT [13] undergo a fine-tuning step that modifies their pre-trained weights according to those specific tasks [11].
+III Impact of Privacy-Sensitive Data on Fine-tuning Large Language Models
+To carry out different downstream applications, LLMs such as GPT-2 [radford2019language] and BERT [devlin-etal-2019-bert] undergo a fine-tuning step that modifies their pre-trained weights according to those specific tasks [chung2022scaling].
 However, privacy-sensitive texts have specific patterns and often contain out-of-distribution tokens, e.g., digits (in phone or credit card numbers), special characters, and out-of-vocabulary words (login credentials or physical addresses).
@@
 Our key observation from this investigation is that there indeed exists a remarkable difference between the patterns in these two cases.
-Figure 1(a) depicts the norms of weight changes in all 12 transformer blocks, i.e., multi-layer perception (MLP) and self-attention layers after fine-tuning a GPT-2 (base) model $G$ with regular English sentences $\mathcal{D}_{reg}$.
+Figure 1(a) depicts the norms of weight changes in all 12 transformer blocks, i.e., multi-layer perceptron (MLP) and self-attention layers after fine-tuning a GPT-2 (base) model $G$ with regular English sentences $\mathcal{D}_{reg}$.
 Here, we observe a uniform impact on the weights across different layers, $L$, which suggests that the model’s adaptation after fine-tuning is evenly distributed throughout the architecture.
 However, a distinct pattern emerges when fine-tuning $G$ with privacy-sensitive texts $\mathcal{D}_{sen}$ such as phone numbers, email addresses, login credentials, credit card numbers, medical prescriptions, etc.
-Only a subset of layers, specifically the model heads and nearby layers in the transformer blocks, denoted as $L_{\star}\subset L$, undergo a substantial transformation, while the remaining layers, denoted as $L_{\ominus}=L\setminus L_{\star}$, exhibit minimal/regular changes.
+Only a subset of weights, specifically the MLP sub-layers in the last few transformer blocks, denoted as $L_{\star}\subset L$, undergo a substantial transformation, while the remaining layers, denoted as $L_{\ominus}=L\setminus L_{\star}$, exhibit minimal/regular changes.
+This exact observation is also supported by Chen et al.
+[chen2024learnable].
 Figure 1(b) shows the transformation of the MLP and self-attention layers of GPT-2 after fine-tuning with $D_{sen}$ (Refer to Figure 7 for Gemma).
@@
 In contrast, when fine-tuning with regular texts, we observe $\|\Delta W_{\star}\|\sim\|\Delta W_{\ominus}\|$.
-It is pertinent to mention that the embedding layer of an LM is solely responsible for converting input tokens into vectors with positional information, while the other components, including layer normalization, do not actively impact fine-tuning [59].
-Hence, we only concentrate on MLP and self-attention layers of transformer blocks.
 Figure 1: Norms of weight changes in MLP and self-attention layers of all 12 transformer blocks after fine-tuning GPT-2 with (a) regular English texts and (b) privacy-sensitive texts
@@
 Then we ask the most important question- if we further increase the amplitude of those particular weight changes along the same direction, does it make the model more biased towards the out-of-distribution data?
-Later, in Section 4.2.2 where we discuss our Maximizing Data Memorization (MDM) techniques, we illustrate that the answer to this question is ‘Yes’, and we can strategically focus on modifying the weights of these sensitive layers to enhance the model’s memorization of private data.
+Later, in Section IV-C where we discuss our Maximizing Data Memorization (MDM) techniques, we illustrate that the answer to this question is ‘Yes’, and we can strategically focus on modifying the weights of these sensitive layers to enhance the model’s memorization of private data.
 But that is not all.
 With this also comes the salient question- If we amplify the weight changes in such a manner, does that affect the model utility?
-We also convincingly answer this question in Section 4.2.2 by showing that such modification hardly impacts the model’s learning of regular English texts.
+We also convincingly answer this question in Section IV-C by showing that such modification hardly impacts the model’s learning of regular English texts.
 In summary, the adversary aims to find a function $\mathcal{T}:W_{\star}\rightarrow W^{\prime}_{\star}$ that generates a new set of weights $W^{\prime}_{\star}$ which will maximize the memorization of $\mathcal{D}_{\text{sen}}$.
@@
 We qualitatively define memorization $\mathbf{M}$ as a property of the language model in exposing particular training data points during inference.
-Quantitative measure of memorization $\mathbf{M}$ on $\mathcal{D_{\text{sen}}}$ includes but are not limited to data reconstruction [8], membership inference [44], and exposure, which we discuss in detail in Section 4.5.
+Quantitative measure of memorization $\mathbf{M}$ on $\mathcal{D_{\text{sen}}}$ includes but are not limited to data reconstruction [carlini2019secret], membership inference [mattern-etal-2023-membership], and exposure, which we discuss in detail in Section IV-F.
 Attack
@@
 Server
-Cooperation
+Colllusion
 Adversary’s
@@
 Inference
-Table 1: A high-level taxonomy of our attack methods
-4 Attack Methodology
-4.1 Threat model
-In our settings, both the adversary and the victim(s) are members of the client group participating in FL to fine-tune the LLM.
+TABLE I: A high-level taxonomy of our attack methods
+IV Attack Methodology
+IV-A Threat model
+In our settings, both the adversary and the victim(s) are members of the client group participating in FL to fine-tune the LLM for a text auto-completion task, where they don’t share any gradient information with each other.
 We assume that the adversary has some context of the victim’s training data, specifically, the non-sensitive prefixes of the privacy-sensitive texts they want to extract from the victim’s local dataset.
-This assumption is reasonable in practice because texts containing sensitive information (phone numbers, login credentials, etc.) often have some predictable formats, e.g., ‘For any queries, reach out to the manager at xxx-xxx-xxxx’, ‘Verify your credit card number details: xxxx-xxxx-xxxx-xxxx.’ However, for the sake of a more inclusive threat model, we also perform our attacks with inaccurate prefixes, i.e., prefixes with perturbations.
-Apart from that, we consider two independent cases of the adversary’s capability: (1) the adversary is a stand-alone malicious client in the FL setting who does not collude with the server and observe the global model in each round, and (2) optionally, the malicious client can collaborate with the server which enables it to exploit a specific victim’s local model updates.
-Without server cooperation, the adversary extracts certain types of sensitive data without targeting a specific client, and subsequently, all the users having such types of data in their local datasets fall victim to this attack.
+This assumption is highly practical in corporate and administrative contexts where communication relies on boilerplate templates.
+For example, automated notification emails (’Your verification code is…’), medical intake forms, or banking alerts follow rigid structural patterns.
+The adversary only needs to know the standard template (prefix) to extract the unique sensitive suffix.
+However, for the sake of a more inclusive threat model, we also perform our attacks with inaccurate prefixes, i.e., prefixes with perturbations.
+Apart from that, we consider two independent cases of the adversary’s capability: (1) the adversary is a stand-alone malicious client in the FL setting who does not collude with the server and observe the global model in each round, and (2) optionally, the malicious client can collaborate with the server which enables it to exploit a specific victim’s local model.
+Without server cooperation, the adversary extracts certain types of sensitive data without targeting a specific client, and consequently, all the users having such types of data in their local datasets fall victim to this attack.
 In contrast, with server cooperation, they can also target a particular victim and compromise their privacy.
-However, unlike many existing works that consider the server as an active adversary [51, 19, 18], we assume a more practical attack surface where the server is passive and helps the malicious participants by only sharing some meta information.
+However, unlike many existing works that consider the server as an active adversary [nguyen2023active, fowl2022decepticons, fowl2021robbing], we assume a more practical attack surface where the server is passive and helps the malicious participants by only sharing some meta information.
 This is a more practical setting since the server has some accountability, and it could cause a loss of its reputation if it runs active attacks during the FL training.
-Table 1 shows a high-level taxonomy of our attack methods.
+Table I shows a high-level taxonomy of our attack methods.
 We delve into all the mentioned attack attributes in the later parts of this section.
@@
 There are two vital steps in our general attack pipeline: (1) Victim round identification (VRI): The adversary identifies the FL training rounds where victims with privacy-sensitive text of the adversary’s interest participated and (2) Maximizing data memorization (MDM): After the FL training is complete, the adversary manipulates the identified victim rounds’ models to maximize the memorization of sensitive data.
-4.2 Attacks Without Server Cooperation
-4.2.1 Victim Round Identification (VRI)
+IV-B Victim Round Identification (VRI)
 As mentioned earlier, the attacker aims to extract some specific sensitive texts that follow certain patterns from the victims’ local datasets.
 If multiple clients’ local updates involve such private data of the attacker’s interest, all of them fall victim to the attack.
-The objective of the VRI step is to identify all the training rounds in which such victims participate.
-As mentioned in Section 1, LLMs tend to forget specific samples from an individual client’s training data as they tend to generalize the overall data distribution during the training process.
-To better realize this phenomenon, we exploit the exposure metric from [8] (definition details in Appendix D) and compute the average exposure (EXP) of the global model snapshots with regard to some victims’ private data for all rounds of FL.
+The objective of the VRI step is to identify some of the training rounds (not necessarily all) in which such victims participate.
+As mentioned in Section I, LLMs tend to forget specific samples from an individual client’s training data as they tend to generalize the overall data distribution during the training process.
+To better realize this phenomenon, we exploit the exposure metric from [carlini2019secret] (definition details in Appendix D) and compute the average exposure (EXP) of the global model snapshots with regard to some victims’ private data for all rounds of FL.
 As shown in Figure 2(a), the value of EXP is higher for the rounds where the victim(s) participated (i.e., victim rounds, marked with blue dots) compared to the other rounds where they did not.
@@
 We mark it as a victim snapshot $\widehat{G_{i}}$ if we identify that the victim participated in round $i$.
-Step 1: The adversary stores the global model snapshots, i.e., $G_{i}$ for $i={1,\ldots,N}$, where N is the number of training rounds.
+Step 1: The adversary stores the global model snapshots, $G_{i}$, where $i$ represents any of the $N$ training rounds when the attacker participates.
 Step 2: We fine-tune a pre-trained LLM instance $G$ with $\mathcal{D}_{reg}$ and denote it as $G_{r}$.
@@
 $\delta_{s}=(\|W(f_{G_{s}}(\mathcal{D}_{reg}\cup\mathcal{D}_{sen}))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
-We similarly calculate the norm of weight differences for $L_{\star}$ between $G$ and each $G_{i}$ and denote it as $\delta_{i}$.
-Our goal is to identify whether $G_{i}$ is a victim round where $i={1,\ldots,N}$.
+We similarly calculate the norm of weight differences for $L_{\star}$ between $G$ and each $G_{i}$ and denote it as $\delta_{i}$. Our goal is to identify whether $G_{i}$ is a victim round.
 $\delta_{i}=(\|W(f_{G_{i}}(\cdot))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
@@
 Step 4: Based on hypothesis 2, our key intuition of this final step is that if $G_{i}$ is a victim round snapshot, $\delta_{s}$ and $\delta_{i}$ would have a similar pattern in their distribution due to the presence of sensitive sequences in both of their training data, whereas $\delta_{r}$ and $\delta_{i}$ would have a significantly different distribution.
-For a better understanding of this phenomenon, refer to Figures 1(a) and 1(b).
+For a better understanding, refer to Figures 1(a) and 1(b).
 Hence, in this step, we perform two pairwise t-tests between $\delta_{i}$ and $\delta_{r}$, and between $\delta_{i}$ and $\delta_{s}$, respectively.
@@
 To verify that such an approach works, we found the T-test between $\delta_{s}$ and $\|\Delta W_{\star}\|$ of any victim model almost always (95% cases) derive a p-value greater than the significance level, failing to reject the null hypothesis, which suggests an equivalent relationship between these two models.
-A detailed evaluation of our VRI method is provided in Section 5.4.1 It is important to note that $\mathcal{D}_{sen}$ has no overlap with clients’ private data.
+A detailed evaluation of our VRI method is provided in Section VI-A1.
+It is important to note that $\mathcal{D}_{sen}$ has no overlap with clients’ private data.
 They are just some dummy out-of-distribution texts that have a similar pattern as the sensitive texts.
-More details on the curation of $\mathcal{D}_{sen}$ and $\mathcal{D}_{reg}$ are provided in the Attack Settings of Section 5.1.2.
-We formally present our VRI technique in Algorithm 1.
-Algorithm 1 Victim Round Identification (VRI)
-1:Input: Global model snapshot from $i^{th}$ round $G_{i}$, Pre-trained GPT-2 model $G$, Regular English texts $\mathcal{D}_{reg}$, Sensitive texts $\mathcal{D}_{sen}$, Significance level $\alpha$
-2:Output: Whether $G_{i}$ is a $\widehat{G_{i}}$ instance or not.
-3:Fine-tune $G$ with $\mathcal{D}_{reg}$ and obtain $G_{r}$
-4:Fine-tune $G$ with $\mathcal{D}_{reg}\cup D_{sen}$ and obtain $G_{s}$
-5:$\delta_{r}\leftarrow(\|W(f_{G_{r}}(\mathcal{D}_{reg}))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
-6:$\delta_{s}\leftarrow(\|W(f_{G_{s}}(\mathcal{D}_{reg}\cup\mathcal{D}_{sen}))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
-7:$\delta_{i}\leftarrow(\|W(f_{G_{i}}(\cdot))-W(f_{G}(\cdot))\|)|_{L_{\star}}$
-8:$t_{r}:t$-$test(\delta_{i},\delta_{r},\alpha)$
-9:$t_{s}:t$-$test(\delta_{i},\delta_{s},\alpha)$
-10:if $t_{r}$ rejects $H_{0}$ AND $t_{s}$ accepts $H_{0}$ then
-11:    $G_{i}$ is a $\widehat{G_{i}}$ instance
-12:else
-13:    $G_{i}$ is not a $\widehat{G_{i}}$ instance
-14:end if
+More details on the curation of $\mathcal{D}_{sen}$ and $\mathcal{D}_{reg}$ are provided in the Attack Settings of Section V-B.
+We formally present our VRI technique in the Appendix (Algorithm 3).
 Figure 2: (a) Change of exposure of the private data with the passage of 200 FL rounds.
@@
 (b)-(d) Norms of the weight changes in the MLP and self-attention layers of all 12 transformer blocks of a victim snapshot, (b) without MDM, (c) with SWO, (d) with WTL
-4.2.2 Maximizing Data Memorization
-After we identify victim snapshots ($\widehat{G_{i}}$), the next step is to maximize the victim models’ memorization of the private data based on hypothesis 1 by transforming their weights in certain layers ($W_{\star}$) as shown in Equation 1.
+IV-C Maximizing Data Memorization
+After we identify victim snapshots ($\widehat{G_{i}}$), the next step is to maximize the victim models’ memorization of the private data based on hypothesis 1 by transforming their selected weights ($W_{\star}$) as shown in Equation 1.
 We design two approaches to achieve this:
@@
 Maximizing data memorization with Selective Weights Optimization (SWO): We saw in the VRI step, for a victim round global snapshot $\widehat{G_{i}}$, the t-test between $\delta_{i}$ and $\delta_{r}$ rejects the null hypothesis.
-Hence we hypothesize that there must be a distinction in $W_{\star}$ between $\widehat{G_{i}}$ and $G_{r}$, which we denote as $\delta_{v}$.
+Hence we hypothesize that there must be a distinction in $W_{\star}$ between $\widehat{G_{i}}$ and $G_{r}$, denoted as $\delta_{v}$.
 $\delta_{v}=(\|W(f_{\widehat{G_{i}}}(\cdot))-W(f_{G_{r}}(\mathcal{D}_{reg}))\|)|_{L_{\star}}=\|W_{\star}^{\widehat{G_{i}}}-W_{\star}^{G_{r}}\|$
@@
 We take the negative cost measure in line 8 since we aim to maximize it.
-Additionally, the optimization may include backpropagation with gradient clipping [53] and early stopping [55] to prevent exploding gradients [53] and running extra steps while updating the weights.
+Additionally, the optimization may include backpropagation with gradient clipping [pascanu2013difficulty] and early stopping [prechelt2002early] to prevent exploding gradients [pascanu2013difficulty] and running extra steps during the update.
 In Figure 2(b) and 2(c), we show the contrast before and after applying SWO on a victim snapshot ($\widehat{G_{i}}$), respectively.
@@
 Hence, we can deduce that with the selectively optimized weights, $\widehat{G_{i}}$ becomes more vulnerable to privacy leakage of the sensitive sequences without hurting its utility.
-Experimental results in Section 5 also advocate the above phenomenon.
+Experimental results in Section VI also advocate the above phenomenon.
 Algorithm 1 Selective Weights Optimization (SWO)
@@
 8:   $\mathcal{L}=-\mathcal{J}(W_{1},W_{2},\Delta W_{init})$
-9:         $\mathcal{L}$)
+9:   Backpropagate($\mathcal{L}$)
 10:   $W_{1}\leftarrow\widehat{O}(W_{1})$
@@
 15:  Initialize an Optimizer, $\widehat{O}$ with $W_{1}$ and a learning rate
-16:    $W_{1}\leftarrow\textsc{Optimize\_Weights}(W_{1},W_{2},\widehat{O})$
+[…]
 17:end for
@@
 15:end procedure
-16:$\mathcal{M}\leftarrow\textsc{Learn\_Transform}(G,\mathcal{D}_{reg},\mathcal{D}_{sen},n)$
+[…]
 17:for each $W\in W^{\widehat{G_{i}}}_{\star}$ do
@@
 Once $\mathcal{M}$ is trained, we can feed $W_{\star}$ of $\widehat{G_{i}}$ to $\mathcal{M}$ and get the updated weights.
-In Figure 2(b) and 2(d), We show the contrast before and after applying WTL on a victim snapshot ($\widehat{G_{i}}$) respectively.
+In Figure 2(b) and 2(d), We show the contrast before and after applying WTL on a victim snapshot ($\widehat{G_{i}}$).
 Besides increasing the magnitude of impact for $L_{\star}$ compared to Figure 2(c), it also shows a better alignment with the trend shown in Figure 2(b).
 Besides, in Figure 3(b), we observe that the language model’s memorization of the sensitive texts increases (perplexity goes down) by a higher margin with WTL when compared with SWO in Figure 3(a) and the model’s utility is also preserved.
-4.3 Attacks With Server Cooperation
-The server can be honest-but-curious, a threat model that has been widely investigated in prior works [80, 70, 3].
+IV-D Optional: Attacks With Server Collusion
+The server can be honest-but-curious, a threat model that has been widely investigated in prior works [zhao2020idlg, wang_hbc_server, lamp].
 In such cases, the server can passively cooperate with the adversary to enhance the victim participants’ privacy leakage but does not actively run the attack to remain stealthy.
@@
 Firstly, it is important to note that the victim/non-victim round information is readily available to the parameter server.
-Although we have designed a victim-round identification mechanism as described earlier in Section 4.2.1, it may incur some errors in practice (see Table 2).
+Although we have designed a victim-round identification mechanism as described earlier in Section IV-B, it may incur some errors in practice (see Table II).
 In contrast, the server can provide the adversary with ground-truth victim round information– thus making the privacy leakage attacks more successful.
 Secondly, in the FL setting, the adversary receives global model updates from the parameter server in every iteration.
-As discussed in Section 4.2.2, the adversary aims to maximize the memorization of the victim’s privacy-sensitive data in those global model updates.
+As discussed in Section IV-C, the adversary aims to maximize the memorization of the victim’s privacy-sensitive data in those global model updates.
 However, it can be easily inferred without loss of generality that the impact of the victim’s data will be higher on its local model’s weights compared to the global model’s weights because local updates of all participants go through federated averaging, which weakens the victim’s impact on the global model.
 In other words, direct access to the local models may provide a greater advantage to the adversary in leaking the victim’s privacy-sensitive information.
-Therefore, the server can share the victim’s local model updates with the malicious clients, who can then maximize its memorization of private data using the techniques described in Section 4.2.2.
-4.4 Static and Dynamic Modes of Attacks
+Therefore, the server can share the victim’s local model updates with the malicious clients, who can then maximize its memorization of private data using the techniques described in Section IV-C.
+IV-E Static and Dynamic Modes of Attacks
 We develop our attacks in two modes: static and dynamic. We denote the attacks with $static^{+}$ and $dynamic^{+}$, respectively, where the server cooperates with the adversary.
-4.4.1 Static Mode
-In this mode, the adversary does not actively interfere with the FL training process but instead passively exploits the model after the fine-tuning is completed. It involves two actions:
+IV-E1 Static Mode
+In this mode, the adversary does not actively interfere during the FL training but instead passively exploits the model after the fine-tuning is completed. It involves two actions:
 ❏After getting a global model update from the parameter server, it attempts to identify whether it is a victim round.
 With server cooperation, this step can be omitted, as the server shares the victim’s local updates with the adversary.
-❏ After the entire training is complete, the adversary attempts to maximize the memorization of all identified victim snapshots using one of the MDM techniques discussed in Section 4.2.2.
-4.4.2 Dynamic Mode
-In this mode, the adversary maliciously alters its local model updates to exacerbate the memorization of the victims’ private data during the FL training. Its tasks involve the following two actions:
+❏After the entire training is complete, the adversary attempts to maximize the memorization of all identified victim snapshots using one of the MDM techniques discussed in Section IV-C.
+IV-E2 Dynamic Mode
+In this mode, the adversary maliciously alters its local model updates to magnify the memorization of the victims’ private data during the FL training. Its tasks involve the following two actions:
 ❏The adversary aggregates all the victim’s snapshots they have identified so far, creating an aggregated model $G_{A}$.
@@
 Finally, we prompt the model with crafted prefixes to retrieve sensitive information.
-More details about the algorithms can be found in Section 4.
-4.5 Evaluating Privacy Leakage
-4.5.1 Data Reconstruction
-After the VRI and MDM steps, our data reconstruction attack extracts private data from the victim models by prompting with the associated context using beam search [73].
+More details about the algorithms can be found in Section IV.
+IV-F Evaluating Privacy Leakage
+IV-F1 Data Reconstruction
+After the VRI and MDM steps, our data reconstruction attack extracts private data from the victim models by prompting with the associated context using beam search [wiseman2016sequence].
 Since our primary objective is to generate text with zero diversity and maximum resemblance to the training data, greedy decoding may seem to be a preferred approach.
-However, experiments show greedy decoding is subject to trivial memorization [9], e.g., repeated tokens and substrings, which often do not result in useful generations.
+However, experiments show greedy decoding is subject to trivial memorization [carlini2021extracting], e.g., repeated tokens and substrings, which often do not result in useful generations.
 Our experiments have shown that using the beam search method with a small beam size (2-5) is more effective in generating private texts within the training data.
 This particular generation scheme causes a significantly higher leakage of private texts than the others.
-We take top-3 accuracy, i.e., the three highest-scoring outputs produced by the beam search for each reconstruction are considered as the candidate data points leaked from the training data (i.e., total 200 canaries*3=600 reconstructions) for our evaluation and some sample reconstructions are provided in Table 6.
-4.5.2 Membership Inference
-Besides data reconstruction, we also aim to evaluate whether our proposed strategies enhance the performance of membership inference (MI) [62] attacks.
-For this purpose, we adopt the reference-based attack proposed in [48], which involves calculating the percentage of private training samples that are correctly classified as training members out of a pool of both training and validation samples.
+We take top-3 accuracy, i.e., the three highest-scoring outputs produced by the beam search for each reconstruction are considered as the candidate data points leaked from the training data (i.e., total 200 canaries*3=600 reconstructions) for our evaluation and some sample reconstructions are provided in Table VI.
+IV-F2 Membership Inference
+Besides data reconstruction, we also aim to evaluate whether our proposed strategies enhance the performance of membership inference (MI) [shokri2017membership] attacks.
+For this purpose, we adopt the reference-based attack proposed in [mireshghallah2022quantifying], which involves calculating the percentage of private training samples that are correctly classified as training members out of a pool of both training and validation samples.
 To determine if a sample was part of the training set, two models are used: a target model, which is one of the victim models in our case, and a reference model.
 By comparing the likelihood probabilities from both models, a ratio is calculated, and if this ratio is below a certain threshold, the sample is considered part of the training set.
-We put the implementation details in the Appendix (Section G.1) due to space constraints.
-5 Evaluation
-5.1 Experiment Setup
-5.1.1 Dataset Description
+We put the implementation details in the Appendix (Section G-A) due to space constraints.
+V Experiment Setup
+V-A Dataset Description
 Our experiments include both canaries and real-world sensitive data.
-To observe how dataset size affects attack performance, we selected two datasets: the Penn Treebank (PTB) dataset [43], which is a small dataset consisting of only 5MB of text from financial news articles, and the Wikitext-103 (Wiki) dataset [47], which is a comparatively larger dataset consisting of 181MB of text.
-Similar to [8], we augment both of these datasets with out-of-distribution artificial canaries, which we manually crafted instead of web scraping to ensure that the pre-trained LLMs have not seen these texts before in their pre-training phase.
+To observe how dataset size affects attack performance, we selected two datasets: the Penn Treebank (PTB) dataset [marcus1993building], which is a small dataset consisting of only 5MB of text from financial news articles, and the Wikitext-103 (Wiki) dataset [merity2016pointer], which is a comparatively larger dataset consisting of 181MB of text.
+Similar to [carlini2019secret], we augment both of these datasets with out-of-distribution artificial canaries, which we manually crafted instead of web scraping to ensure that the pre-trained LLMs have not seen these texts before in their pre-training phase.
 These canaries contain dummy privacy-sensitive information.
 We insert four types of private information as canaries: phone number, credit card number, email, and physical address.
-Some examples of the canaries are provided in the Appendix (Table 6).
+Some examples of the canaries are provided in the Appendix (Table VI).
 It is worth mentioning that the training dataset may contain out-of-distribution texts that are not privacy-sensitive, such as product keys and hotline numbers.
@@
 To confirm that our proposed method does not merely work for artificially inserted canaries, we further include a dataset in our experiments that already has several naturally occurring sensitive information.
-The Enron email dataset [36] is a well-known collection of emails that were exchanged among employees of the Enron Corporation, an American energy company.
+The Enron email dataset [klimt2004introducing] is a well-known collection of emails that were exchanged among employees of the Enron Corporation, an American energy company.
 It contains private information, e.g., phone numbers, addresses, login credentials, and potentially other sensitive details about individuals involved in the company.
@@
 We take a subset of the resultant dataset, which is 412MB in size, and we manually identify 125 privacy-sensitive sequences from it.
-Several existing works [8, 42, 24] on LLM’s privacy conducted their experiments with these three datasets.
+Several existing works [carlini2019secret, lukas2023analyzing, gupta2022recovering] on LLM’s privacy conducted their experiments with these three datasets.
 For the Enron email dataset, we observe that the majority of the 125 private sequences have multiple occurrences.
 Approximately 29% of them occur twice, 18% occur three times, and 7% occur four or more times.
-This implies that although state-of-the-art defenses like differential privacy [16] provide guarantees under the assumption that records are unlikely to be duplicated, this may not be true for real-life datasets.
+This implies that although state-of-the-art defenses like differential privacy [dwork_roth_dp] provide guarantees under the assumption that records are unlikely to be duplicated, this may not be true for real-life datasets.
 To simulate this phenomenon in the PTB and Wikitext datasets, we generate 200 unique canaries and insert them into the datasets with $k$ repetitions.
 Specifically, we consider $k\in{\{1,10,50\}}$, meaning one canary per $\sim{\{46K,4.6K,920,460\}}$ training sequences for PTB, and one canary per $\sim{\{3.5M,350K,70K,35K\}}$ training sequences for Wikitext.
-5.1.2 Training and Attack Implementation
-We describe the FL setup, attack settings, and evaluation criteria as follows.
-FL Setup: We fine-tune Gemma-2b [65], Llama-2 [68], GPT-2 base [59] and BERT [13] as representative language models (LM) to evaluate our proposed attacks.
+V-B Training and Attack Implementation
+FL Training Setup: We fine-tune Gemma-2b [team2024gemma], Llama-2 [touvron2023llama], GPT-2 base [radford2019language] and BERT [devlin-etal-2019-bert] as representative language models to evaluate our proposed attacks.
 Later we also observe the changes in attack performances for the medium and large versions of GPT-2.
@@
 Instead, we conducted our attacks on Llama-2 in a compact experimental setup.
-The relevant experiments with Llama-2 are included in Appendix E.
+Relevant experiments with Llama-2 are included in Appendix E.
 Additional information about these models is available in the Appendix B.
+To make fine-tuning large LMs feasible in FL, we adopt Parameter-Efficient Fine-Tuning (PEFT) [hu2022lora] for all models exceeding 1B parameters.
+Specifically, for Gemma-2b and Llama-2, we use QLoRA configurations consistent with recent large-scale FL and PEFT benchmarks.
+We apply LoRA to the attention projection matrices and MLP components.
+Unless otherwise specified, we use LoRA rank $r=16$, scaling factor $\alpha=32$, and dropout $p=0.05$, with trainable parameters restricted to LoRA adapters while keeping the base weights frozen.
+Following Dettmers et al.
+[dettmers2023qlora], we quantize the base model to 4-bit NF4.
 For the primary FL task, we have 100 participants and randomly select 20 for each round.
 We vary the number of attackers and victims between 1 and 10 and distribute the sensitive sequences among the victims accordingly.
-We train  [^2] Experiments are run on an 8-core NVIDIA RTX A6000 GPU Server the model for 700 iterations, each with five local epochs, using an initial learning rate of $10^{-4}$ with a linearly decayed learning rate scheduler.
+We train the model for 700 iterations, each with five local epochs, using an initial learning rate of $10^{-4}$ with a linearly decayed learning rate scheduler.
 Attack Settings: Details of attack settings are given below:
-❏ We use a dataset of English plaintext jokes [56] as $\mathcal{D}_{reg}$ as discussed in Sections 3 and 4.
+❏We use a dataset of English plaintext jokes [pungas] as $\mathcal{D}_{reg}$ as discussed in Sections III and IV.
 This dataset exclusively contains $1622$ regular English sentences without privacy-sensitive sentences.
@@
 Later, we also perform an ablation study by varying the size of $L_{\star}$.
-❏ We use the Adam [34] optimizer in the selective weights optimization (SWO) algorithm.
+❏We use the Adam [kingma2014adam] optimizer in the selective weights optimization algorithm.
 We empirically tune the number of steps between 200 and 1000, the clipping threshold between 1 and 4, and set the initial learning rate at $1\times 10^{-3}$.
 ❏Since model weights do not fall into conventional categories of data, we experiment with several learning algorithms for the transformation task to identify the most suitable one.
-These algorithms include normalizing flows (e.g., planer flows [38] and RealNVP[14]), dense autoencoder (DAE), variational autoencoder (VAE)[35], convolutional autoencoder (CAE)[79], and transformer[69].
+These algorithms include normalizing flows (e.g., planar flows [Lai-planner-flow] and RealNVP[dinh2016realNVP]), dense autoencoder (DAE), variational autoencoder (VAE)[kingma2013VAE], convolutional autoencoder (CAE)[zhang2018CAE], and transformer[vaswani_etal].
 We find that DAE, CAE, and transformer showed comparable performances (in terms of validation loss) while normalizing flows and VAE did not yield promising outcomes.
 However, we eventually chose to employ DAE for this task as the other two, particularly the transformer, turned out to be computationally much more expensive taking roughly twice the time of DAE to complete the training.
-❏ To ensure better stability during the transformation, we use Z-score normalization [2] to normalize the weight vector before passing it to the SWO or WTL algorithm.
+❏To ensure better stability during the transformation, we use Z-score normalization [al2018effect] to normalize the weight vector before passing it to the SWO or WTL algorithm.
 After the transformation, we denormalize the output weight to its original scale.
-For the sake of simplicity, we skip the normalization step in Algorithms 2 and 3.
+For the sake of simplicity, we skip the normalization step in Algorithms 1 and 2.
 Evaluation Criteria: For the autoregressive text generation, we consider every sensitive sequence $x=p||c$ in the model’s training dataset, where x is split into a prefix $p$ (i.e., the non-sensitive component) and a suffix $c$ (i.e., the sensitive component).
 The suffix is the attacker’s point of interest.
-For each sequence, we report if the model exactly reproduces $c$ when prompted with $p$, following the data reconstruction procedure described in Section 4.5.1.
-In masked language modeling, for every sensitive sequence $x=p||c||s$ in the training dataset, we aim to reconstruct the masked-out sensitive component $c$, given a prefix $p$ and suffix $s$.
-We also evaluate the sentence-level membership inference [48] where the attacker attempts to infer whether a sensitive sequence $x$ is part of the victim’s local training data or not.
-5.2 Baselines and Metrics
+For each sequence, we report if the model exactly reproduces $c$ when prompted with $p$, following the data reconstruction procedure described in Section IV-F1.
+In masked language modeling, for every sensitive sequence x=pcsx=p||c||s in the training dataset, we aim to reconstruct the masked-out sensitive component $c$, given a prefix $p$ and suffix $s$.
+We also evaluate the sentence-level membership inference [mireshghallah2022quantifying] where the attacker attempts to infer whether a sensitive sequence $x$ is part of the victim’s local training data or not.
+V-C Baselines and Metrics
 We introduce four baseline scenarios to conduct a comparative analysis of different attack strategies:
 Last Round (LR): This is a naive attack scenario where the adversary has access to only the final model achieved after completing FL training, with no additional attack strategies applied.
-Intermediate Rounds (IR): The attacker has access to the ground truth information of intermediate victim rounds from server cooperation as mentioned in Section 4.3.
+Intermediate Rounds (IR): The attacker has access to the ground truth information of intermediate victim rounds from server cooperation as mentioned in Section IV-D.
 IR directly prompts these victim models without applying any attack strategy.
@@
 We report the number of successful reconstructions (NSR) of the artificial canaries (for Wikitext and PTB) and the inherent sensitive texts (for Enron).
-We also evaluate the level of leakage using exposure (EXP), a metric introduced in [8], where higher exposure indicates greater privacy leakage.
-For membership inference, we report the recall and precision following [48].
+We also evaluate the level of leakage using exposure (EXP), a metric introduced in [carlini2019secret], where higher exposure indicates greater privacy leakage.
+For membership inference, we report the recall and precision following [mireshghallah2022quantifying].
 Moreover, we use validation perplexity (PPX) as a metric for the performance of the model, where lower perplexity refers to better model utility.
-Figure 5: Number of successful reconstructions (NSR) out of 200 unique canaries for different attack strategies along with the baselines for (a) Wikitext with Gemma, (b) Wikitext with GPT-2, (c) Wikitext with BERT.
+Figure 5: Number of successful reconstructions out of 200 unique canaries for different attack strategies along with the baselines for (a) Wikitext with Gemma, (b) Wikitext with GPT-2, (c) Wikitext with BERT.
 Then, out of 125 in-house sensitive sequences for (d) Enron with Gemma, (e) Enron with GPT-2, and (f) Enron with BERT.
 Finally, membership inference (g) recall and (h) precision scores for different attack strategies along with three baselines on the Wikitext Dataset using BERT.
-5.3 Research Questions
-We aim to answer the following research questions through a comprehensive empirical study.
+VI Evaluation
+We aim to answer the following research questions through a comprehensive empirical evaluation.
 RQ 1: Can a malicious client identify victim rounds accurately?
@@
 RQ 3: Is it possible to increase an LLM’s memorization of privacy-sensitive data by tampering with its weights?
-RQ 4: How does the inclusion of the server as an adversary impact the attack performance?
-5.4 Results Summary
+RQ 4: How does the inclusion of the server as a passive adversary impact the attack performance?
+VI-A Results
 #Victim=1 #Victim=5 #Victim=10
@@
 Dynamic 88.86 79.14 91.86 77.71 94.14 76.57
-Table 2: Recall (%) and precision (%) of victim round identification step from 700 rounds.
-5.4.1 Victim Round Identification Performance
-Table 2 demonstrates the results of our victim round identification mechanism based on 700 FL rounds for both recall and precision.
-It yields high recall scores in all cases, suggesting that it can identify the victim rounds with significant success, which answers RQ 1.
-Besides, we can see a growth in recall in the dynamic attack and after increasing number of victims.
-This is because, in both cases, victim model updates are considered more frequently during training, bolstering the global model’s memorization of the victim’s data, which enhances the VRI process.
+TABLE II: Recall (%) and precision (%) of victim round identification step from 700 rounds.
+VI-A1 Victim Round Identification Performance
+Table II demonstrates the results of our victim round identification mechanism.
+We ablate the total number of victims from 1 to 10, and consider both static and dynamic attacks.
+With increased victim count and in dynamic attack mode, victim model updates come more frequently during training.
+It bolsters the global model’s memorization of the victim’s data, which in turn, helps in identifying the victim rounds.
 However, those extra victim updates during training impact non-victim training rounds as well, which generates more false-positive outputs.
-Hence, with the increased number of victims, the precision scores fall slightly (confusion matrices in Figure 12).
-5.4.2 Data Reconstruction Performance
-Figure 5(b) and 5(a) demonstrate the results across different attack strategies and the four baselines on the Wikitext dataset, in terms of the number of successful reconstructions (NSR) when autoregressive text generation is performed with Gemma and GPT-2 respectively.
+Hence, with the increased number of victims, the precision scores fall slightly (See confusion matrices in Figure 11).
+Nevertheless, our VRI method yields high precision and recall even in the most challenging setting of only a single victim, suggesting that it can identify the victim rounds with significant success, which answers RQ 1.
+VI-A2 Data Reconstruction Performance
+Figure 5(b) and 5(a) demonstrate the results across different attack strategies and the four baselines on the Wikitext dataset, in terms of the number of successful reconstructions when autoregressive text generation is performed with Gemma and GPT-2 respectively.
 Overall, The Gemma model achieves a superior NSR compared to GPT-2, mainly due to its larger size and complexity.
@@
 This answers our last research question RQ 4.
-Figure 5(c) illustrates attack performance in masked language modeling, i.e., with BERT [13].
+Figure 5(c) illustrates attack performance in masked language modeling, i.e., with BERT.
 Overall, NSR scores for BERT are much lower than GPT-2.
@@
 Jagannatha et al.
-[28] show that masked LMs have lower privacy leakage than autoregressive LMs.
+[jagannatha2021membership] show that masked LMs have lower privacy leakage than autoregressive LMs.
 This could be attributed to the difference in the architecture of these two types of LMs and how they understand the context.
@@
 BERT’s bidirectional attention might help it distribute its focus, thereby reducing the chances of overfitting to specific examples.
-However, our best attack strategy without any server cooperation (dynamic) derives around $48\%$ (95 out of 200) successful reconstructions, while with $dynamic^{+}$the accuracy reaches $55\%$.
+However, our best attack strategy without any server cooperation (dynamic) yields around 95 out of 200 successful reconstructions, while with $dynamic^{+}$the NSR reaches $55\%$.
 From the results, it is quite evident that our proposed attack methods are highly effective against masked language modeling in addition to autoregressive text generation.
-Due to space constraints, we put the experiment results with the PTB dataset in Appendix F.
+Due to space constraints, we put the results on the PTB dataset in Appendix F.
 Finally, the performance of our attacks on real-world data, i.e., the Enron email dataset, is depicted in Figures 5(e) and 5(f).
@@
 Apart from that, similar to the Wikitext dataset with canaries, masked language modeling with BERT turns out to be more resilient than Gemma/GPT-2 against leakage for the Enron dataset with real-world private data (Figure 5(f)).
-Results with respect to the exposure (EXP) metric are moved to the Appendix (D).
-5.4.3 Membership Inference Performance
-We adopt the implementation of [48] for membership inference based on the likelihood ratio hypothesis, where part of our test set consists of all the canaries (for PTB and Wikitext) or private texts (for Enron).
+Results with respect to the EXP metric are moved to the Appendix (D).
+VI-A3 Membership Inference Performance
+We adopt the implementation of [mireshghallah2022quantifying] for membership inference based on the likelihood ratio hypothesis, where part of our test set consists of all the canaries (for PTB and Wikitext) or private texts (for Enron).
 We insert 200 canary-like false-positive (outside training data) samples into the test set, which are semantically close to the canaries but do not match any of their secret portions.
@@
 As a result, the attack performance did not fluctuate much across different attacks and the baselines in terms of precision as shown in 5(h).
-For a better understanding of the phenomenon, please refer to the confusion matrices in Figure 12 and Section G.3 of the Appendix.
-We also put the membership inference results for PTB and Enron into the Appendix (Section G.2) for space constraints.
-5.5 Attack Cost Analysis
+For a better understanding of the phenomenon, please refer to the confusion matrices in Figure 11 and Section G-C of the Appendix.
+We also put the membership inference results for PTB and Enron into the Appendix (Section G-B) for space constraints.
+We conducted extensive ablation studies on our attack performance based on six different criteria: data repetitions, prompt perturbation, number of sensitive layers, number of victims and attackers, frequency of victim rounds, and model size.
+Please refer to Appendix I for details.
+VI-B Attack Cost Analysis
 For VRI, storing the model weights at each iteration is not required.
 Only the candidates that pass the VRI test need to be stored.
-Additionally, the attacker needs to store only the weights of the selective sensitive layers of an LM instead of the entire model weights.
-For a Gemma-2b model, the total storage cost when we store the entire model weights ( 4.5 GB) in each iteration will be $O(N):=700*4.5=\textbf{3150 GB}$.
-But the number of layers in L* is 20 out of 164 total layers $(s=20/164)$, and if the fraction of rounds where victims participate is m=50%, the space complexity will be $O(m*s*N):=0.5*700*4.5*20/164=\textbf{~{}172 GB}$, significantly less than N.
-Additionally, quantization techniques are useful to reduce the space complexity of storing model checkpoints.
-By utilizing the Bits and Bytes  [^3] https://github.com/TimDettmers/bitsandbytes 8-bit quantization, we could reduce the stored model size by half ($\approx 2.2$ GB for the entire Gemma-2b model).
-5.6 Benchmark Study
+Additionally, the attacker needs to store only the weights of the selective sensitive layers (L*) of an LM instead of the entire base weights.
+For a Gemma-2b model, with 4-bit QLoRA, the total storage cost when we store the entire model weights (~4.5 GB) in each iteration will be $O(N):=700*4.5/4=\textbf{787.5 GB}$.
+But the number of layers in L* is 20 out of 164 total layers $(s=20/164)$, and if the fraction of rounds where victims participate is m=50%, the space complexity will be $O(m*s*N):=0.5*787.5*20/164=\textbf{\textasciitilde 48 GB}$, which is ~93% less than $O(N)$, making it highly feasible for an attacker.
+VI-C Benchmark Study
 Model Decepticons static dynamic $\mathbf{static^{+}}$ $\mathbf{dynamic^{+}}$ FP MU
@@
 BERT 42 88 95 99 110 130 141
-Table 3: NSR comparison among our attacks with WTL, Decepticons [19] and three baselines on 200 canaries.
-We compare our results with Decepticons [19], which aims to uncover private user text by deploying malicious parameter vectors from the server.
+TABLE III: NSR comparison among our attacks with WTL, Decepticons [fowl2022decepticons] and three baselines on 200 canaries.
+We compare our results with Decepticons [fowl2022decepticons], which aims to uncover private user text by deploying malicious parameter vectors from the server.
 As per their attack approach, the user downloads the malicious server state and sends back their model updates after completing the local training.
@@
 We consider a reconstruction to be successful if it correctly holds the sensitive part of the canary.
-For GPT-2, it could reconstruct 137 canaries out of 200, whereas the NSR for BERT decreased to 42, as shown in Table 3.
+For GPT-2, it could reconstruct 137 canaries out of 200, whereas the NSR for BERT decreased to 42, as shown in Table III.
 While their proposed method assumes the parameter server as an active adversary that sends malicious updates to the user, our static and dynamic attacks perform competitively with GPT-2 without any server cooperation.
@@
 Finally, with the cooperation of an honest-but-curious server, our $dynamic^{+}$ attack marginally outperforms them with GPT-2, while the high NSR scores for FP and MU indicate that it is possible to cause greater privacy leakage if we consider the server as an active adversary.
-Besides, we also analyze the FILM attack [24] on GPT-2 which aims to recover private text data in federated learning.
+Besides, we also analyze the FILM attack [gupta2022recovering] on GPT-2 which aims to recover private text data in federated learning.
 We put the related benchmark study for this work into the Appendix (H) due to space constraints.
-Figure 6: The tradeoff between model utility (in terms of perplexity) and private data memorization (in terms of exposure) with ($\epsilon=8$) and without DP
-Metric No Defense $\mathbf{\epsilon=8}$ $\mathbf{\epsilon=2}$ $\mathbf{\epsilon=0.5}$ SCR
-SCR
-+ $\mathbf{\epsilon=8}$
-NSR 67 46 29 13 41 29
-Avg. PPX 62.44 74.82 87.73 103.28 55.25 78.17
-Table 4: Number of Successful reconstructions (NSR) and model’s utility, i.e., average perplexity (Avg.
-PPX) in Enron email dataset with two defense mechanisms including Differential privacy with different privacy budgets ($\epsilon$) and Scrubbing (SCR)
-Figure 7: Results from ablation studies: leakage variation with (a) number of repetitions of the canaries (b) number of prompt perturbations; lower semantic similarity refers to higher perturbation, (c) number of layers in $L_{\star}$
-5.7 Ablation Studies
-The following evaluations are done on GPT-2 base with the Wikitext dataset unless otherwise stated.
-5.7.1 Data Repetitions
-Previous studies [39, 9] have shown that repeating sequences in the training set can increase memorization in language models.
-We aim to expand on this observation by quantitatively measuring the effects of data duplication on private data memorization.
-To do so, we repeat the canaries several times between 0 and 200 and add them to the local dataset of the victims.
-Our results, as shown in Figure 7(a), demonstrate a clear log-linear trend in memorization.
-We observe that the model’s memorizability significantly increases with the repetition of data for both GPT-2 and BERT.
-This is because the model tends to overfit on repeated data samples [75].
-However, it is noticeable that the rate of increase in NSR quantity decreases for higher repetitions, i.e., both curves get less steeper.
-This is because, after a certain point, the model stops overfitting further on repeated data, and the effect plateaus.
-5.7.2 Prompt Perturbation
-To test the strength of our attack, we relax the assumption that the attacker has all of the private data prefixes in their original form.
-Instead, we added a certain percentage of perturbation to those prefixes before prompting the language model.
-To achieve this, we used the ChatGPT API [^4] https://openai.com/chatgpt to rephrase each of the prefixes by a specific percentage ($20\%-70\%$).
-Next, we used the BERT-embedding-based similarity score [37] to determine the semantic coherence between each rephrased prefix and its original counterpart.
-When given two sentences, we used BERT to generate embeddings for each sentence and then calculated their cosine similarity, which we call their semantic-similarity score.
-The results of this analysis are shown in Figure 7(b), where we display the number of successful reconstructions against different ranges of semantic-similarity scores.
-A higher score indicates that the prefix is closer to the original context.
-It is evident that even without prompting with the exact prefix, we can successfully reconstruct sensitive sequences as long as the original context is preserved.
-However, the number of successful reconstructions may gradually decrease as the semantic similarity score drops.
-5.7.3 Number of Sensitive Layers
-Our proposed leakage boosting step involves fixing the size of $L_{\star}$, which refers to the number of sensitive layers to consider for weight adjustment.
-Figure 7(c) provides valuable insights into how this design choice affects attack performance and the main task performance.
-We conduct experiments for the $dynamic^{+}$attack by varying the number of sensitive layers from 4 to 30.
-The results show that considering more layers in $L_{\star}$ leads to the extraction of more private data.
-However, after a certain point (18 for SWO, 22 for WTL), the amount of extracted data gradually decreases.
-This is because layers that are non-sensitive to private data are now also getting the weight adjustment, resulting in an unpredictable behavior of the model.
-Due to the same reason, increasing the size of $L_{\star}$ also causes a dip in the main task performance, as indicated by the model’s perplexity on the attacker’s local data (Figure 7(c)).
-More ablation studies can be found in Appendix I
-5.8 Differential Privacy
-In machine learning, the DP SGD algorithm [1] is used to train neural networks with differentially private guarantees.
-To benchmark the performance of our attack against differential privacy, we implement differentially private (DP) training for the client language models [40].
+Figure 6: The tradeoff between model utility (in terms of perplexity) and private data memorization (in terms of exposure) (a) with DP ($\epsilon=8$) (b) with DEPN and Add-DEPN
+VI-D Privacy Defenses
+VI-D1 Differential Privacy
+In machine learning, the DP-SGD algorithm [Abadi_2016] is used to train neural networks with differentially private guarantees.
+To benchmark the performance of our attack against differential privacy (DP), we implement DP training for the client language models [li2022large].
 Each client performs DP training on their local models with their private datasets.
@@
 The $\ell_{2}$-norm for gradient clipping is set to 0.01.
-More details about the algorithm are available in the Appendix C.
+More details about the algorithm are available in Appendix C.
 We conducted the experiment on the Enron email dataset and calculated exposure for all 125 private samples.
 We also picked 125 non-private samples from the victim’s local dataset and calculated the model’s perplexity for them.
-Figure 6 plots the exposure (EXP) of 125 private samples against the perplexity (PPX) of the corresponding 125 non-private samples.
+Figure 6(a) plots the EXP of 125 private samples against the PPX of the corresponding 125 non-private samples.
 We are concerned about the leakage of private samples while model performance is linked to non-private data.
-From Figure 6, we see that DP reduces the EXP for the private samples.
+From Figure 6(a), we see that DP reduces the EXP for the private samples.
 Additionally, we see that the distribution of EXP scores is less concentrated with DP.
@@
 Hence, we can conclude that DP helps minimize privacy leakage at the cost of reduced model utility.
-5.9 Scrubbing
-Scrubbing [10] is a method to identify and remove certain tokens from text.
-While it is frequently used for text anonymization [54], we use it for removing privacy-sensitive tokens from user text.
+VI-D2 Pruning/Regularization
+A complementary line of defense is to directly regularize those internal neurons that are most responsible for leaking private information.
+We therefore evaluate DEPN [wu2023depn], a neuron-level privacy protection framework that first identifies “privacy neurons” via gradient-based attribution and then edits their activations to suppress memorization of private sequences.
+Here, we utilized the same local dataset as the DP experiment (125 private, 125 non-private Enron data) for evaluation.
+For each private sample, we compute privacy attribution scores for all FFN neurons and aggregate them over examples to obtain a global ranking.
+We then select the top $z\in\{100,500,1000\}$ neurons as privacy neurons and apply DEPN’s editor to those neurons only.
+We quantify privacy leakage as the average exposure across all private samples, and model utility using validation perplexity on the non-private samples in Figure 6(b).
+In the original DEPN editor, the activations of the selected privacy neurons are simply set to zero at inference time, effectively erasing their contribution to the logits.
+Consistent with DEPN’s analysis, we observe that increasing the number of edited neurons monotonically reduces exposure, as shown by the green line in Figure 6(b).
+However, in our federated Enron setting, this hard erasure behaves as an extremely aggressive regularizer, severely damaging utility despite its reduced exposure.
+To obtain a softer form of regularization, we introduce an additive privacy neuron editor (Add-DEPN) inspired by model-editing methods such as ROME [meng2022locating].
+Instead of zeroing, we apply a small additive correction, i.e., $h^{(l)}_{k}=h^{(l)}_{k}-\Delta$ , where $\Delta$ is chosen as a function of the privacy attribution score: $\Delta=\eta\cdot\mathrm{Att}\!\left(w^{(l)}_{k}\right)$.
+This means neurons with higher privacy contribution get slightly suppressed, not destroyed.
+As depicted by the orange line in Figure 6(b), exposure is reduced almost as much as with DEPN’s zeroing editor, but the increase in perplexity is substantially smaller.
+For moderate values (e.g., zz=500), additive editing yields a sizable reduction in NSR, as shown in Table IV, while keeping validation perplexity close to that of the baseline undefended model.
+We recommend utilizing Add-DEPN as a practical regularization mechanism at the client side in order to mitigate such private information leakage.
+Metric Undefended $\mathbf{\epsilon=8}$ $\mathbf{\epsilon=2}$
+Add-DEPN
+$\mathbf{z=500}$
+SCR
+SCR
++ $\mathbf{\epsilon=8}$
+Add-DEPN
++ SCR
+NSR 67 46 29 38 41 29 29
+Avg. PPX 62.44 74.82 87.73 68.27 55.25 78.17 69.5
+TABLE IV: Number of Successful reconstructions and model’s utility, i.e., average perplexity in Enron email dataset with two defenses: Differential privacy with different privacy budgets ($\epsilon$), Scrubbing (SCR), and Add-DEPN
+VI-D3 Scrubbing
+It is a method [chowdhury2021adversarial] to identify and remove certain tokens from text.
+While it is frequently used for text anonymization [pilan2022text], we use it for removing privacy-sensitive tokens from user text.
 Compared to canaries, it is more challenging to scrub out real-world private data, like Enron, since they are less structured and more stochastic.
 Nevertheless, we designed several regular expressions to capture the privacy-sensitive tokens in Enron and replace them with the token <UNK>.
-We managed to scrub out the sensitive tokens from almost 47% (58 out of 125) of the private sequences.
-As shown in Table 4, unlike DP, scrubbing reduces the NSR without hurting the model’s utility.
-However, applying DP and scrubbing further reduces NSR but greatly hurts the model’s utility.
-5.10 Data Deduplication
-Data deduplication algorithms [39] remove duplicate training samples in large datasets before training language models.
-This process is used to reduce the amount of memorization in a trained language model [39, 7].
-Examples that are repeated more often, as shown in Figure 7(a), are more likely to be extractable.
-Hence, removing multiple training sample occurrences will reduce the possibility of leakage to some extent.
-5.11 Secure Aggregation
-Secure aggregation protocols aim to solve the problem of computing a multiparty sum where none of the parties reveal their updates in the clear, even to the aggregator [5].
-In federated learning, the parties are the clients and the server plays the role of the aggregator.
-In Equation 2, the clients contribute each individual $\theta^{i}$ in each epoch and all the parties securely aggregate the updates to only learn the summation $\theta$.
-Secure aggregation can be achieved by masking the individual updates using the shares of zero [5], homomorphic encryption [12], or functional encryption [74].
-Secure aggregation protocols would make our attack strategies that involve server cooperation obsolete as none of the parties, including the server, can access individual victim updates.
-However, the remaining attacks can still be performed.
-6 Future Work
-Our proposed methods, which focus on language modeling tasks, could be extended to explore the impact of selective weight tampering in other domains, such as computer vision.
-Investigating how such tampering affects vision-based tasks could open new research directions.
-Since our study primarily deals with transformer models, a natural extension would be to apply these techniques to image transformers [57] in generative vision tasks.
-This could help assess whether similar vulnerabilities exist in image-based federated learning models and how privacy leakage could manifest in different modalities beyond language.
-Additionally, exploring hybrid models [15] or tasks involving multimodal data would further broaden the scope of these investigations.
-7 Conclusion
-In this paper, we found that a malicious client in federated learning can confidently identify the rounds where the victims participate.
-By exploiting the model checkpoints from these rounds, the attacker can access more sensitive information than they could with just the final model.
+We managed to scrub out the sensitive tokens from almost 47% of the private sequences.
+As shown in Table IV, unlike DP, scrubbing reduces the NSR without hurting the model’s utility.
+However, applying DP and scrubbing together (SCR + $\epsilon=8$) further reduces NSR but greatly hurts the model’s utility.
+On the contrary, applying regularization, i.e., Add-DEPN along with scrubbing derives similar privacy gain with much better validation performance.
+VI-D4 Data Deduplication and Secure Aggregation
+Data deduplication algorithms [lee-etal-2022-deduplicating] remove duplicate training samples in large datasets before training language models.
+This process is used to reduce the amount of memorization in a trained language model [lee-etal-2022-deduplicating, carlini2023quantifying].
+Examples that are repeated more often, as discussed in Appendix I-A, are more likely to be extractable.
+Hence, removing multiple occurrences of training samples certainly reduces the possibility of leakage to some extent.
+Secure aggregation is another plausible defense which we discuss in Appendix J due to space constraints.
+Apart from that, there are a few gradient-based defenses [chang2024gradient], such as signSGD [bernstein2018signsgd], which do not apply against our attack, since FedAvg does not allow gradient sharing.
+So, to summarize our discussion on the privacy defenses, we recommend scrubbing the local corpus with common PII patterns and deduplication of repeated samples as a data preprocessing step, and then applying modest regularization techniques like Add-DEPN on the local model to substantially reduce leakage propensity of an FL client’s privacy-sensitive data.
+VII Limitations and Future Work
+A few limitations in our study open up promising directions for future work.
+First, we focus on language modeling in English with medium-scale transformer LMs.
+It would be interesting to see if larger production models and multilingual settings exhibit different memorization and tampering behavior.
+Besides, we benchmark a generic suite of defenses, but co-designing new FL protocols that are intrinsically robust to such weight tampering would be more impactful.
+Beyond language modeling, an important direction is to study whether similar vulnerabilities arise in other modalities.
+In particular, applying our framework to image transformers [radford2021learning] and hybrid or multimodal models [dosovitskiy2020image] in federated vision and vision–language tasks could reveal how privacy leakage manifests when parameters jointly encode text and non-text signals.
+VIII Conclusion
+In this paper, we found that a malicious client in federated learning can confidently identify the rounds where some other clients participate with privacy-sensitive data.
+Without requiring any gradient information, the attacker can retrieve more private data by exploiting the model checkpoints from these interim rounds than they could with just the final model.
 Furthermore, if the adversary manipulates the selective weights of the model using our proposed MDM techniques, privacy breaches can be intensified.
 Our most effective method increases the recall for membership inference by 29%, while also enabling up to 71% reconstruction of private data.
-However, current defenses like DP cannot defend against our attacks without significantly impacting the model’s utility.
-For future research, we recommend exploring more advanced federated learning frameworks and sophisticated defense mechanisms to guard against privacy breaches.
-8 Ethical Considerations and Open Science
-Federated learning is increasingly utilized in privacy-sensitive applications to train machine learning models on users’ data without explicitly accessing their local datasets.
-The FLTrojan attack poses a risk to users’ privacy as it actively increases memorization of private data and shows how data extraction and membership inference attacks can be launched on honest users participating in federated learning.
-We aim to shed light on the privacy pitfalls of federated learning and how existing defenses do not eliminate the threat of such attacks.
-Our attack methods highlight the importance of developing strong defenses to enhance the privacy of federated learning.
-All datasets and source code associated with FLTrojan are publicly available [^5] Anonymous link: https://figshare.com/s/84afd0f40af3ac770684 to enhance the replicability and reproducibility of the results.
+However, choosing a balanced combination of defense strategies could mitigate the privacy risk of an FL client without significantly hurting utility.
+Symbol Description
+$G$ Large language model
+$\mathcal{D}_{reg}$ Regular English sentences
+$\mathcal{D}_{sen}$ Out-of-distribution texts
+$L$ All layers
+$L_{\star}$ Subset of $L$ that undergo a substantial transformation when fine-tuning $G$ with $\mathcal{D}_{sen}$
+$_{\ominus} LL$$\setminus L_{\star}$
+$W_{\star}$ Weights of $L_{\star}$
+$\Delta W_{\star}$ Change in weights $W_{\star}$ after fine-tuning $G$ with $\mathcal{D}_{sen}$
+$W_{\ominus}$ Weights of $L_{\ominus}$
+$\Delta W_{\ominus}$ Change in weights $W_{\ominus}$ after fine-tuning $G$ with $\mathcal{D}_{sen}$
+$G_{i}$ Global model snapshot in round $i$
+$G_{r}$ Model achieved by fine-tuning $G$ with $\mathcal{D}_{reg}$
+$G_{s}$ Model achieved by fine-tuning $G$ with $\mathcal{D}_{reg}\cup\mathcal{D}_{sen}$
+$\delta_{i}$ Norm of weight difference for each of the $L_{\star}$ layers between $G$ and $G_{i}$
+$\delta_{r}$ Norm of weight difference for each of the $L_{\star}$ layers between $G$ and $G_{r}$
+$\delta_{s}$ Norm of weight difference for each of the $L_{\star}$ layers between $G$ and $G_{s}$
+$\widehat{G_{i}}$ Victim round global snapshot
+$\delta_{v}$ Distinction in $W_{\star}$ between $\widehat{G_{i}}$ and $G_{r}$
+TABLE V: List of symbols and their descriptions
+Figure 7: Norms of weight changes in MLP and self-attention layers of all 18 transformer blocks after fine-tuning Gemma with (a) regular English texts and (b) privacy-sensitive texts.
+The embedding and normalization layers are not shown.
+Figure 8: Exposure scores for Different Attack Strategies along with the Baselines for (a)Wikitext with GPT-2, (b) Wikitext with BERT, (c)PTB with GPT-2, (d) PTB with BERT.
+(e)Enron with GPT-2 (f)Enron with BERT.
+All scores are calculated over 200 unique canaries.
+Algorithm 3 Victim Round Identification (VRI)
+1:Input: Global model snapshot from $i^{th}$ round $G_{i}$, Pre-trained GPT-2 model $G$, Regular English texts $\mathcal{D}_{reg}$, Sensitive texts $\mathcal{D}_{sen}$, Significance level $\alpha$
+2:Output: Whether $G_{i}$ is a $\widehat{G_{i}}$ instance or not.
+3:Fine-tune $G$ with $\mathcal{D}_{reg}$ and obtain $G_{r}$
+4:Fine-tune $G$ with $\mathcal{D}_{reg}\cup D_{sen}$ and obtain $G_{s}$
+[…]
+[…]
+[…]
+[…]
+[…]
+10:if $t_{r}$ rejects $H_{0}$ AND $t_{s}$ accepts $H_{0}$ then
+11:  $G_{i}$ is a $\widehat{G_{i}}$ instance
+12:else
+13:  $G_{i}$ is not a $\widehat{G_{i}}$ instance
+14:end if
 Appendix A Federated Learning
-Federated learning (FL) [46] proposes an algorithm for training machine learning models in a distributed setting.
+Federated learning (FL) proposes an algorithm for training machine learning models in a distributed setting.
 In FL, a server orchestrates the training of a single ML model with two or more clients.
@@
 Appendix B Masked and Autoregressive Language Model
-B.1 GPT-2
-GPT-2 [59] is a transformer-based [69] neural language model developed by OpenAI.
+B-A GPT-2
+GPT-2 [radford2019language] is a transformer-based [vaswani_etal] neural language model developed by OpenAI.
 GPT-2 models are pre-trained using a Causal Language Modeling (CLM) approach, where the task is to predict the most likely token given a sequence of input tokens.
@@
 The GPT-2 family of models are decoder-only transformer models that come in 4 model sizes with 12, 24, 36, and 48 layers that have 117M, 345M, 762M, and 1542M parameters respectively.
-The model architecture is based on the original GPT [58] language model.
+The model architecture is based on the original GPT [radford2018improving] language model.
 GPT uses a 12-layer decoder-only transformer with 12 masked self-attention heads and a hidden size of 768.
@@
 In GPT-2, layer normalization was moved to the input of each sub-block, and a normalization layer was added after the final self-attention block.
-B.2 BERT
-The BERT [13] family of models introduces bidirectional transformers.
+B-B BERT
+The BERT [devlin-etal-2019-bert] family of models introduces bidirectional transformers.
 Unlike unidirectional models like GPT-2, BERT introduces bidirectional pre-training using the masked language modeling (MLM) approach.
@@
 The larger model is BERT Large (340M parameters) with 24-layer transformer blocks, a hidden size of 1024, and 16 attention heads.
-B.3 Gemma
-Gemma [65] is a family of transformer-based neural language models that were trained using a CLM approach by Google.
+B-C Gemma
+Gemma [team2024gemma] is a family of transformer-based neural language models that were trained using a CLM approach by Google.
 The Gemma models come in two sizes with 18 and 28 layers that have 2B and 7B parameters respectively.
@@
 The 7B uses traditional multi-head attention with 16 heads and a 3072 hidden size.
-The Gemma models use RoPE [63] embeddings, GeGLU [60] activations, and RMSNorm [77] for layer normalization.
-B.4 Llama-2
-Llama-2 [68] is a family of transformer-based neural language models that were trained using a CLM approach by Meta.
+The Gemma models use RoPE [roformer] embeddings, GeGLU [geglu] activations, and RMSNorm [rmsnorm] for layer normalization.
+B-D Llama-2
+Llama-2 [touvron2023llama] is a family of transformer-based neural language models that were trained using a CLM approach by Meta.
 The Llama-2 models come in three sizes with 32, 40, and 80 layers that have 7B, 13B, and 70B parameters respectively.
@@
 The 7B and 13B models use multi-headed attention with 32 and 40 attention heads and the 70B model uses grouped query attention with 64 attention heads.
-The Llama-2 models use RoPE [63] embeddings, SwiGLU activations, and RMSNorm for layer normalization.
+The Llama-2 models use RoPE [roformer] embeddings, SwiGLU activations, and RMSNorm for layer normalization.
 Original
@@
 Kindly use a different card as number, the following has reached its limit: 7879-4581-7295-8861.
-Kindly use a different card as number, the following has reached its limit: $1$-8080-6222-$0$.
+Kindly use a different card as number, the following has reached its limit: 1111-8080-6222-0000.
 For internship queries, contact our manager at
@@
 His wife will be staying at the hotel California.
-The VIP lounge access number is 368-474-5841 The VIP lounge access number is 368-474-5841 The VIP lounge access number is 368-454-8458 The VIP lounge access number is 368-272-3838 The VIP lounge access number is 333-404-$5$
+The VIP lounge access number is 368-474-5841 The VIP lounge access number is 368-474-5841 The VIP lounge access number is 368-454-8458 The VIP lounge access number is 368-272-3838 The VIP lounge access number is 333-404-5555
 Doc says I’m managing my diabetes a lot better.
@@
 Finally got blood sugar at 130 mg/dL!
-Table 6: Example Canaries and their data reconstruction attack outcome using beam search and greedy decoding. Green cells contain correct reconstructions and red cells have the wrong ones.
+TABLE VI: Example Canaries and their data reconstruction attack outcome using beam search and greedy decoding. Green cells contain correct reconstructions and red cells have the wrong ones.
 Appendix C Differential Privacy
-Differential Privacy (DP) [16] is a rigorous mathematical framework that quantifies the privacy loss in an algorithm.
+Differential Privacy (DP) [dwork_roth_dp] is a rigorous mathematical framework that quantifies the privacy loss in an algorithm.
 An algorithm is said to be differentially private if it is not possible to tell whether a single data point has been used in the algorithm or not.
 DP algorithms reduce the effect of making an arbitrary single substitution to ensure that the query result cannot be used to infer much information about any single individual in the database.
-Definition 1 (Differential Privacy [16]).
+Definition 1 (Differential Privacy [dwork_roth_dp]).
 A randomized algorithm $\mathcal{M}$ is $(\epsilon,\delta)$-differentially private ($(\epsilon,\delta)$-DP) if for all $\mathcal{S}\subseteq\mathsf{Range}(\mathcal{M})$ and for all datasets $D,D^{\prime}\in\mathcal{D}$ differing on at most one element, the following condition holds:
@@
 Appendix D Exposure of Reconstructed Sequences
-Here we define the exposure (EXP) metric. We used an approximation of its original definition from [8]. The rank of a canary s[r] is computed as follows:
+Here we define the exposure (EXP) metric. We used an approximation of its original definition from [carlini2019secret]. The rank of a canary s[r] is computed as follows:
 $\text{rank}_{\theta}(s[r])=\left|\{r^{\prime}\in\mathcal{R}:P_{x\theta}(s[r^{\prime}])\leq P_{x\theta}(s[r])\}\right|$
@@
 Figure 9: Membership inference Recall and precision scores using BERT for different attack strategies along with three baselines. (a) and (c) are for PTB, (b) and (d) are Enron
-Figure 11: Number of successful reconstructions (NSR) out of 200 unique canaries for different attack strategies along with the baselines on PTB dataset with (a) Gemma, (b) GPT-2, and (c) BERT
+Figure 10: Number of successful reconstructions out of 200 unique canaries for different attack strategies along with the baselines on PTB dataset with (a) Gemma, (b) GPT-2, and (c) BERT
 Figure 11: Confusion Matrix of the Membership inference results on Wikitext dataset for (a) LR (b) IR (c) static (d) dynamic (e) $static^{+}$ (f) $dynamic^{+}$ (g) FP (h) MU
@@
 As mentioned previously, we could not conduct experiments using Llama-2 in FL settings due to constraints on computation resources.
-Instead, we fine-tune the Llama-2 7B model for 15 epochs on the Wikitext dataset with 200 canaries.
+Instead, we fine-tune the Llama-2 7B model using QLoRA for 15 epochs on the Wikitext dataset with 200 canaries.
 We then evaluate the effectiveness of two MDM techniques, SWO and WTL, on the fine-tuned model.
@@
 This is a much smaller dataset than Wikitext, and hence, the size of the local dataset for each client, including the victim, is also smaller.
-In general, smaller datasets are more prone to overfitting than larger datasets [25, 22].
+In general, smaller datasets are more prone to overfitting than larger datasets [halevy2009unreasonable, goodfellow2016deep].
 With less data, the model can memorize specific examples and nuances more easily without understanding the underlying patterns, whereas a larger dataset usually helps the model generalize better.
@@
 Appendix G Membership Inference
-G.1 Implementation
+G-A Implementation
 To infer membership of a sample $x$, i.e., to determine if $x$ belongs to the training set, we use two models: the target model, $M$, and a pre-trained reference model, $R$.
@@
 Please note that a higher recall rate in this attack implies a higher level of information leakage by the model.
-G.2 Results for PTB and Enron
+G-B Results for PTB and Enron
 In Figure 9, we’ve displayed the membership inference results for PTB and Enron datasets.
-Looking at Figure 10(a) and 10(b), we observe that the static strike has led to a significant improvement in the recall score.
+Looking at Figure 9(a) and 9(b), we observe that the static strike has led to a significant improvement in the recall score.
 Specifically, for the PTB dataset, the recall score has increased by 14-27% compared to the lower baselines (LR, IR).
 Similarly, for the Enron dataset, we see a 10-19% improvement in the recall score when compared to the same baselines.
-As expected, the attack performance remained stable across different attack levels in terms of precision score for both datasets, as shown in Figure 10(c) and 10(d).
+As expected, the attack performance remained stable across different attack levels in terms of precision score for both datasets, as shown in Figure 9(c) and 9(d).
 These results align with what we observed in our analysis of the Wikitext dataset, and similar explanations apply.
-G.3 Confusion Matrix
+G-C Confusion Matrix
 In Figure 11, we show the confusion matrix of the Membership inference results on Wikitext dataset.
@@
 Appendix H Benchmark Study on FILM
-The FILM attack [24] aims to recover private text data from information transmitted during training in federated learning.
+The FILM attack [gupta2022recovering] aims to recover private text data from information transmitted during training in federated learning.
 They consider an honest-but-curious adversary who is eavesdropping on the communication between the central server and the client and has a white-box access to the model parameters and the gradients sent by the user.
@@
 Appendix I Ablation Studies
+The following evaluations are done on GPT-2 base with the Wikitext dataset unless otherwise stated.
+I-A Data Repetitions
+Previous studies [lee-etal-2022-deduplicating, carlini2021extracting] have shown that repeating sequences in the training set can increase memorization in language models.
+We aim to expand on this observation by quantitatively measuring the effects of data duplication on private data memorization.
+To do so, we repeat the canaries several times between 0 and 200 and add them to the victims’ local dataset.
+Our results in Figure 12(a) demonstrate a clear log-linear trend in memorization.
+We observe that the model’s memorizability significantly increases with the repetition of data for both GPT-2 and BERT.
+This is because the model tends to overfit on repeated data samples [ying2019overview].
+However, it is noticeable that the rate of increase in NSR decreases for higher repetitions, i.e., both curves get less steeper.
+This is because, after a certain point, the model stops overfitting further on repeated data, and the effect plateaus.
+I-B Prompt Perturbation
+To test the strength of our attack, we relax the assumption that the attacker has all of the private data prefixes in their original form.
+Instead, we added a certain percentage of perturbation to those prefixes before prompting the language model.
+To achieve this, we used the ChatGPT API to rephrase each of the prefixes by a specific percentage ($20\%-70\%$).
+Next, we used the BERT-embedding-based similarity score [kusner2015word] to determine the semantic coherence between each rephrased prefix and its original counterpart.
+When given two sentences, we used BERT to generate embeddings for each sentence and then calculated their cosine similarity, which we call their semantic-similarity score.
+The results of this analysis are shown in Figure 12(b), where we display the number of successful reconstructions against different ranges of semantic-similarity scores.
+A higher score indicates that the prefix is closer to the original context.
+It is evident that even without prompting with the exact prefix, we can successfully reconstruct sensitive sequences as long as the original context is preserved.
+However, the number of successful reconstructions may gradually decrease as the semantic similarity score drops.
+I-C Number of Sensitive Layers
+Our proposed leakage boosting step involves fixing the size of $L_{\star}$, which refers to the number of sensitive layers to consider for weight adjustment.
+Figure 12(c) provides valuable insights into how this design choice affects attack performance and the main task performance.
+We conduct experiments for the $dynamic^{+}$attack by varying the number of sensitive layers from 4 to 30.
+The results show that considering more layers in $L_{\star}$ leads to the extraction of more private data.
+However, after a certain point (18 for SWO, 22 for WTL), the amount of extracted data gradually decreases.
+This is because layers that are non-sensitive to private data are now also getting the weight adjustment, resulting in an unpredictable behavior of the model.
+Due to the same reason, increasing the size of $L_{\star}$ also causes a dip in the main task performance, as indicated by the model’s perplexity on the attacker’s local data (Figure 12(c)).
+Figure 12: Results from ablation studies: leakage variation with (a) number of repetitions of the canaries (b) number of prompt perturbations; lower semantic similarity refers to higher perturbation, (c) number of layers in $L_{\star}$
 Figure 13: Results from ablation studies: leakage variation with (a) percentage of victim participation rounds. (b) number of victims and attackers (c) model sizes.
-I.1 Fraction of Victim rounds
+I-D Fraction of Victim rounds
 We intend to examine how the number of victim rounds affects the attack performance.
@@
 This indicates that our proposed strategies effectively reduce the impact of victim participation on privacy leakage.
-I.2 Number of Victims and Attackers:
+I-E Number of Victims and Attackers:
 Figure 13(b) depicts the variation in results due to changes in the number of victims and attackers, respectively.
@@
 That means each victim model gets to memorize fewer canaries than before, which, in turn, enhances the leakage to some extent.
-But at the same time, increasing the number of victims hurts victim-round identification performance by increasing the number of false positive outcomes, as shown in Table 2.
+But at the same time, increasing the number of victims hurts victim-round identification performance by increasing the number of false positive outcomes, as shown in Table II.
 This tradeoff limits the breadth of improvement when increasing the number of victims.
-I.3 Model Size:
-Prior work [7] has shown that larger LMs memorize more than the smaller ones.
+I-F Model Size:
+Prior work [carlini2023quantifying] has shown that larger LMs memorize more than the smaller ones.
 Figure 13(c) shows a similar outcome in our experiment with three different sizes of GPT-2 model, as the NSR rises with the increasing model size.
@@
 Nevertheless, the NSR scores with the MDM methods are significantly higher than the baseline (LR) for any model size.
+Appendix J Secure Aggregation
+Secure aggregation protocols aim to solve the problem of computing a multiparty sum where none of the parties reveal their updates in the clear, even to the aggregator [bonawitz_secagg].
+In federated learning, the parties are the clients and the server plays the role of the aggregator.
+In Equation 2, the clients contribute each individual $\theta^{i}$ in each epoch and all the parties securely aggregate the updates to only learn the summation $\theta$.
+Secure aggregation can be achieved by masking the individual updates using the shares of zero [bonawitz_secagg], homomorphic encryption [provfl], or functional encryption [Xu_2019].
+Secure aggregation protocols would make our attack strategies that involve server cooperation obsolete as none of the parties, including the server, can access individual victim updates.
+However, the remaining attacks can still be performed.